### v1

# On a mean-field Pontryagin minimum principle for stochastic optimal control

Abstract.
This papers outlines a novel extension of the classical Pontryagin minimum (maximum) principle to stochastic optimal control problems. Contrary to the well-known stochastic Pontryagin minimum principle involving forward-backward stochastic differential equations, the proposed formulation is deterministic and of mean-field type. The Hamiltonian structure of the proposed Pontryagin minimum principle is achieved via the introduction of an appropriate gauge variable. The gauge freedom can be used to decouple the forward and reverse time equations; hence simplifying the solution of the underlying boundary value problem. We also consider infinite horizon discounted cost optimal control problems. In this case, the mean-field formulation allows converting the computation of the desired optimal control law into solving a pair of forward mean-field ordinary differential equations. The proposed mean-field formulation of the Pontryagin minimum principle is tested numerically for a controlled inverted pendulum and a controlled Lorenz-63 system.
1. Introduction
We consider the optimal control problem for a controlled stochastic differential equation (SDE) of the form
(1) ${\rm d}X_{t}=b(X_{t}){\rm d}t+G(X_{t})U_{t}{\rm d}t+\Sigma^{1/2}{\rm d}B_{t},\qquad X_{0}=a,$
under either the finite horizon cost function
(2) $J_{T}(a,U_{0:T})=\mathbb{E}\left[\int_{0}^{T}\left(c(X_{t})+\frac{1}{2}U_{t}^{\rm T}R^{-1}U_{t}\right){\rm d}t+f(X_{T})\right]$
or the infinite horizon discounted cost function
(3) $J(a,U_{0:\infty})=\mathbb{E}\left[\int_{0}^{\infty}e^{-\gamma t}\left(c(X_{t})+\frac{1}{2}U_{t}^{\rm T}R^{-1}U_{t}\right){\rm d}t\right].$
Here $B_{t}$ denotes $d_{x}$-dimensional Brownian motion, $\Sigma\in\mathbb{R}^{d_{x}\times d_{x}}$ the symmetric positive definite diffusion matrix, $R\in\mathbb{R}^{d_{u}\times d_{u}}$ a symmetric positive definite weight matrix, $G(x)\in\mathbb{R}^{d_{x}\times d_{u}}$ the possibly position-dependent control matrix, $c(x)\geq 0$ the running cost, $f(x)\geq 0$ the terminal cost, and $\gamma>0$ the discount factor. See, for example, [2] for more details. We also introduce the weighted norm $\|\cdot\|_{R}$ via
(4) $\|u\|^{2}_{R}=u^{\rm T}R^{-1}u.$
The aim is to find the closed loop control law $U_{t}=u_{t}(X_{t})$ that minimizes the cost $J_{T}$ over the set of admissible control laws.
It is well known, that the associated optimal control problems for ordinary differential equations (ODEs); i.e. $\Sigma=0$ in (1), can be solved by the Pontryagin minimum (maximum) principle (PMP) [21]. The PMP has been extended to SDEs (1) in terms of forward-backward SDEs [20, 2]. In this paper, we provide an alternative extension via a deterministic mean-field formulation, which leads to interacting particle implementations in the form of Hamiltonian ODEs. The formulation is attractive from a computational perspective since it allows for a decoupling of the forward and backward evolution equations for finite horizon optimal control problems and a forward simulation for infinite horizon optimal control problems. The proposed mean-field approach also delivers closed loop control laws in contrast to the classical PMP for ODEs [21]. The proposed methodology can then also be used to approximate viscosity solutions for controlled ODEs by setting $\Sigma=\epsilon I$ for sufficiently small $\epsilon$ [3].
Mean-field formulations and their interacting particle approximations have been considered before in the context of the Hamilton–Jacobi–Bellman (HJB) equation [2]. The key idea is to connect the time evolution of the value function to a probability density function evolving under an appropriate nonlinear Fokker–Planck equation and its mean-field representation [14, 10, 25, 11]. In this paper, we directly approximate the gradient of the value function and derive appropriate Hamiltonian evolution equations; thus extending mean-field formulations to the PMP.
The remainder of the paper is structured as follows. The following Section 2 summarizes key mathematical results for finite horizon optimal control problems at the level of the underlying HJB and Fokker–Planck equations [2, 19]. The novel mean-field formulation of the PMP is provided in Section 3 starting from an Eulerian variational perspective in terms of the HJB equation. Following [27, 9], the arising action functional is subsequently transformed into a Lagrangian formulation with trajectories following mean-field Hamiltonian ODEs. The infinite horizon optimal control problem is treated in Section 4. Here the key result is a pair of forward mean-field ODEs, which asymptotically deliver the desired closed loop control law. A number of numerical approximation techniques are discussed in Section 5; including an approximation based on discrete Schrödinger bridges [16, 28] as outlined in Section 5.2. The mean-field formulation is tested for two infinite horizon discounted cost (3) optimal control problems in Section 6. The first example is provided by a controlled inverted pendulum [18] while the second example is provided by a controlled Lorenz-63 system [12].
2. Mathematical background
It is well-known [3, 2] that, assuming sufficient regularity, the desired closed loop control law is provided by
(5) $u_{t}(x)=-RG(x)^{\rm T}\nabla_{x}v_{t}(x)$
with the optimal value function $v_{t}(x)$ satisfying the HJB equation
(6) $-\partial_{t}v_{t}=b\cdot\nabla_{x}v_{t}+\frac{1}{2}\Sigma:D_{x}^{2}v_{t}+c-\frac{1}{2}\|RG(x)^{\rm T}\nabla_{x}v_{t}\|^{2}_{R},\qquad v_{T}=f.$
Here $A:B=\sum_{i,j}a_{ij}b_{ij}$ [19]. In case of deterministic dynamics; i.e. $\Sigma=0$ in (1), the PMP provides the open loop optimal control $U_{t}=-RG(X_{t})^{\rm T}P_{t}$ along the solution $X_{t}$ of (1) by solving the ODEs
(7a) $\dot{X}_{t}$ $=b(X_{t})-G(X_{t})RG(X_{t})^{\rm T}P_{t},$
(7b) $-\dot{P}_{t}$ $=(D_{x}b(X_{t}))^{\rm T}P_{t}-\frac{1}{2}\nabla_{x}\|RG^{\rm T}(X_{t})P_{t}\|_{R}^{2}+\nabla_{x}c(X_{t})$
in $(X_{t},P_{t})$ subject to the boundary conditions $X_{0}=a$ and $P_{T}=\nabla_{x}f(X_{T})$ [21]. Formulation (7) offers significant computational savings compared to (6) and has been widely used especially in the context of nonlinear model predictive control (NMPC) [7, 23]. The PMP for ODEs has been extended to SDEs (1) giving rise to the, so called, stochastic PMP with (7) replaced by a pair of forward-backward SDEs [2].
In this paper, we develop a Hamiltonian mean-field formulation in the spirit of the PMP-based Hamiltonian ODE formulation (7), which, in turn, can be approximated by interacting particle systems. Since the PMP directly targets the gradient of the value function $v_{t}(x)$; i.e., $P_{t}=y_{t}(X_{t})$ with $y_{t}(x)=\nabla_{x}v_{t}(x)$, we note that
(8) $-\partial_{t}y_{t}=\nabla_{x}(y_{t}\cdot b)+\frac{1}{2}\nabla_{x}(\Sigma:D_{x}y_{t})+\nabla_{x}c-\frac{1}{2}\nabla_{x}\|RG^{\rm T}y_{t}\|^{2}_{R},\quad y_{T}=\nabla_{x}f,$
which follows form (6) assuming sufficient regularity of the value function.
Substituting the optimal control law (5) into (1), the law $\pi_{t}$ of the states $X_{t}$ satisfies the Fokker–Plank equation
(9) $\partial\pi_{t}=-\nabla_{x}\cdot\left(\pi_{t}(b+Gu_{t})\right)+\frac{1}{2}\nabla_{x}\cdot(\Sigma\nabla_{x}\pi_{t}),\qquad\pi_{0}=p_{0},$
for given initial density $p_{0}$ [19].
We assume throughout this paper that the diffusion matrix $\Sigma$ is position independent. However, the proposed PMP extends to position dependent $\Sigma(x)$. We also allow for $\Sigma=0$, in which case our formulation provides the optimal closed loop control law instead of the open loop solution provided by (7). In this case, adding diffusion via $\Sigma=\epsilon I$ with $\epsilon\to 0$ sufficiently small allows one to numerically approximate viscosity solutions of (6) [3].
3. Pontryagin mean-field formulation
In this section, we develop a novel mean-field approach to the PMP starting from an Eulerian perspective in the spirit of ideal fluid dynamics [27]. Specifically, we combine the HJB equation (6) and a Liouville equation for the flow of densities $\rho_{t}(x)$ under a single variational principle. In order to do so, we assume that the initial states $X_{0}$ are distributed according to a given distribution $p_{0}(x)$. Following [27], we then transform the variational principle into a Lagrangian framework in terms of evolving positions $X_{t}\in\mathbb{R}^{d_{x}}$ and momenta $P_{t}\in\mathbb{R}^{d_{x}}$. These equations are Hamiltonian and provide the novel PMP formulation. While we consider the finite horizon optimal control problem (2) in this section; an extension to infinite horizon optimal control problems can be found in Section 4.
3.1. Eulerian formulation
In this subsection, we develop a variational formulation for the HJB equation (6). The first step is to formally decouple the diffusion term in (6) through an additional function $\phi_{t}$, i.e.,
(10a) $-\partial_{t}v_{t}$ $=b\cdot\nabla_{x}v_{t}+\frac{1}{2}\Sigma:D_{x}^{2}\phi_{t}+c-\frac{1}{2}\|RG(x)^{\rm T}\nabla_{x}v_{t}\|^{2}_{R},$
(10b) 00 0 $=\phi_{t}-v_{t}.$
The second step is to introduce Lagrange multipliers $\rho_{t}(x)$ and $\beta_{t}(x)$ in order to enforce (10) in terms of the action functional
(11a) $\mathcal{S}(v,\rho,\beta)$ $=-\int_{0}^{T}\int_{\mathbb{R}^{d_{x}}}\left\{\partial_{t}v_{t}+\nabla_{x}v_{t}\cdot b+c-\frac{1}{2}\|RG^{\rm T}\nabla_{x}v_{t}\|^{2}_{R}\right\}\rho_{t}{\rm d}x\,{\rm d}t\,+$
(11b) $\qquad\qquad\int_{0}^{T}\int_{\mathbb{R}^{d_{x}}}\left\{\beta_{t}\cdot\nabla_{x}(\phi_{t}-v_{t})+\frac{1}{2}\nabla_{x}\cdot(\Sigma\nabla_{x}\phi_{t})\,\right\}\rho_{t}{\rm d}x\,{\rm d}t\,+$
(11c) $\qquad\qquad\int_{\mathbb{R}^{d_{x}}}\left\{(v_{T}-f)\,\rho_{T}-v_{0}p_{0}\right\}{\rm d}x,$
where we have used $\Sigma:D_{x}^{2}v=\nabla_{x}\cdot(\Sigma\nabla_{x}v)$.
More specifically, the Lagrange multiplier $\beta_{t}$ enforces the constraint (10b) (up to a constant), while the Lagrange multiplier $\rho_{t}$ enforces the HJB formulation (10a). Furthermore, variation with respect to $v_{t}$ leads, after integration by parts, to the Liouville equation
(12) $\partial_{t}\rho_{t}=-\nabla_{x}\cdot\left(\rho_{t}(b-\beta_{t}-GRG^{\rm T}\nabla_{x}v_{t})\right),\qquad\rho_{0}=p_{0},$
in the density $\rho_{t}$. The associated ODE is given by
(13) $\dot{X}_{t}=b(X_{t})-\beta_{t}(X_{t})-G(X_{t})RG(X_{t})^{\rm T}\nabla_{x}v_{t}(X_{t}),\qquad X_{0}\sim p_{0}.$
While (10a) is obtained independently of the choice of $\beta_{t}$, a further variation with respect to $\phi_{t}$ leads to
(14) $-\nabla_{x}\cdot(\rho_{t}\beta_{t})+\frac{1}{2}\nabla_{x}\cdot(\Sigma\nabla_{x}\rho_{t})=0$
and, hence, to
(15) $\beta_{t}=\frac{1}{2}\Sigma\nabla_{x}\log\rho_{t}.$
With this choice of $\beta_{t}$, the associated Liouville equation (12) turns into the Fokker–Planck equation (9) for the SDE (1). In other words, (13) and (15) lead to an ODE mean-field reformulation of the SDE (1) as considered previously, for example, in [15].
3.2. Lagrangian formulation
Let us now convert (11) into a corresponding Lagrangian formulation [27]. Let $X_{\tau}(a)$ denotes the solution of (13) at time $\tau=t$ with initial condition $X_{0}=a$. Here we have introduced $\tau$ as a new time variable in order to distinguish the two time derivatives arising from the Lagrangian (total time-derivative) and Eulerian (partial time-derivative) perspective. Furthermore, the initial conditions $a\sim p_{0}$ may be viewed as labels in the sense of Lagrangian fluid dynamics. We therefore consider functions $x(a)$, $p(x)$, $u(a)$, and $\beta(a)$ and note that
(16) $\int_{\mathbb{R}^{d_{x}}}g(X_{\tau}(a))\,p_{0}(a){\rm d}a=\int_{\mathbb{R}^{d_{x}}}g(x)\,\rho_{t}(x){\rm d}x$
for any function $g(x)$. Next we convert (11) into its corresponding Lagrangian formulation and consider in particular
(17a) $\mathcal{I}$ $:=-\int_{0}^{T}\int_{\mathbb{R}^{d_{x}}}\partial_{t}v_{t}(x)\,\rho_{t}(x){\rm d}x\,{\rm d}t$
(17b) $=-\int_{0}^{T}\int_{\mathbb{R}^{d_{x}}}\partial_{t}v_{t}(X_{\tau=t}(a))\,p_{0}(a){\rm d}a\,{\rm d}t$
(17c) $=\int_{0}^{T}\int_{\mathbb{R}^{d_{x}}}P_{\tau}(a)\cdot\dot{X}_{\tau}(a)\,p_{0}(a){\rm d}a\,{\rm d}\tau-\int_{\mathbb{R}^{d_{x}}}\left(v_{T}(X_{T}(a))-v_{0}(X_{0}(a))\right)\,p_{0}(a){\rm d}a,$
where we have introduced the momentum variable
(18) $P_{\tau}(a)=\nabla_{x}v_{\tau}(X_{\tau}(a)),$
which amounts to an irrotational fluid in the sense of [27]. Here we have used
(19) $\partial_{t}v_{t}(x=X_{\tau=t}(a))=\partial_{\tau}v_{\tau}(X_{\tau}(a))-\nabla_{x}v_{\tau}(X_{\tau}(a))\cdot\dot{X}_{\tau}(a)$
as well as the fact that the total time derivative $\partial_{\tau}v_{\tau}(X_{\tau}(a))$ contributes boundary terms only. Furthermore,
(20a) $\int_{\mathbb{R}^{d_{x}}}\beta_{t}(x)\cdot\nabla_{x}(\phi_{t}(x)-v_{t}(x))\,\rho_{t}(x){\rm d}x=$
(20b) $\qquad\qquad\int_{\mathbb{R}^{d_{x}}}\tilde{\beta}_{\tau}(a)\cdot\nabla_{x}\left(\phi_{\tau}(X_{\tau}(a))-X_{\tau}(a)\cdot P_{\tau}(a)\right)\,p_{0}(a){\rm d}a$
with $\tilde{\beta}_{\tau}(a)=\beta_{t}(X_{\tau}(a))$.
The equations of motion in $(X_{\tau},P_{\tau})$ arise now from the phase space action principle [4, 9] applied to
(21) $\mathcal{S}(X,P,\beta)=\int_{\mathbb{R}^{d_{x}}}\left\{\int_{0}^{T}\left(P_{\tau}\cdot\dot{X}_{\tau}-H(X_{\tau},P_{\tau},\beta_{\tau})\right){\rm d}\tau-f(X_{T})\right\}p_{0}{\rm d}a$
with Hamiltonian density
(22a) $H(x,p,\beta)(a)$ $:=-\frac{1}{2}\|RG(x(a))^{\rm T}p(a)\|^{2}_{R}+p(a)^{\rm T}b(x(a))+c(x(a))\,+$
(22b) $\qquad\qquad\beta(a)\cdot\nabla_{x}\left(\phi(x(a))-x(a)^{\rm T}p(a)\right)+\frac{1}{2}\Sigma:D_{x}^{2}\phi(x(a)).$
The desired optimal closed loop control law (5) is provided by
(23) $u_{\tau}(X_{\tau}(a))=-RG(X_{\tau}(a))^{\rm T}P_{\tau}(a).$
Upon introducing the generator
(24) $\mathcal{L}_{\beta}g=\beta\cdot\nabla_{x}g+\frac{1}{2}\Sigma:D_{x}^{2}g$
and
(25) $g(x)=\phi(x)-x^{\rm T}p,$
we can rewrite (22) in more compact form as
(26a) $H(x,p,\beta)(a)$ $:=-\frac{1}{2}\|RG(x(a))^{\rm T}p(a)\|^{2}_{R}+p(a)^{\rm T}b(x(a))+c(x(a))\,+$
(26b) $\qquad\qquad\mathcal{L}_{\beta}\left(\phi(x(a))-x(a)^{\rm T}p(a)\right).$
Variation with respect to $\beta_{\tau}$ leads to the constraint
(27) $\nabla_{x}\phi_{\tau}(X_{\tau}(a))=P_{\tau}(a),$
which defines the function $\phi_{\tau}(x)$ in terms of $X_{\tau}(a)$ and $P_{\tau}(a)$ up to a constant. Using the thus specified gradient of $\phi_{\tau}(x)$, we obtain the closed loop control
(28) $u_{\tau}(x)=-RG(x)^{\rm T}\nabla_{x}\phi_{\tau}(x).$
Finally, variations with respect to $X_{\tau}$ and $P_{\tau}$ lead to the Hamiltonian evolution equations
(29a) $\dot{X}_{\tau}(a)$ $=\nabla_{p}H(X_{\tau}(a),P_{\tau}(a),\beta_{\tau}(a)),$
(29b) $-\dot{P}_{\tau}(a)$ $=\nabla_{x}H(X_{\tau}(a),P_{\tau}(a),\beta_{\tau}(a))$
for each $a\in\mathbb{R}^{d_{x}}$. The boundary conditions are $X_{0}(a)=a\sim p_{0}$ and $P_{T}(a)=\nabla_{x}f(X_{T}(a))$. The equations (29) are of mean-field type through $\phi_{t}(x)$, which is determined by the constraint (27).
Dropping labels $a\in\mathbb{R}^{d_{x}}$ from now on and denoting time by $t$ again, the Hamiltonian equations of motion (29) become
(30a) $\dot{X}_{t}$ $=b(X_{t})-G(X_{t})RG(X_{t})^{\rm T}P_{t}-\beta_{t}$
(30b) $=b(X_{t})-G(X_{t})RG(X_{t})^{\rm T}P_{t}-\mathcal{L}_{\beta_{t}}X_{t},$
(30c) $-\dot{P}_{t}$ $=\left(D_{x}b(X_{t})\right)^{\rm T}P_{t}\,+$
(30d) $\qquad\qquad\nabla_{x}\left\{c(X_{t}^{(i)})-\frac{1}{2}\|RG(X_{t})^{\rm T}P_{t}\|^{2}_{R}+\mathcal{L}_{\beta_{t}}(\phi_{t}(X_{t})-X_{t}\cdot P_{t})\right\}$
(30e) $=\left(D_{x}b(X_{t})\right)^{\rm T}P_{t}-\frac{1}{2}\nabla_{x}\|RG(X_{t})^{\rm T}P_{t}\|^{2}_{R}+\nabla_{x}c(X_{t}^{(i)})+\mathcal{L}_{\beta_{t}}\nabla_{x}\phi_{t}(X_{t}),$
which are to be compared to the classical PMP formulation (7). The only new terms arise in connection with the generator $\mathcal{L}_{\beta_{t}}$.
The following theorem provides the key result with regard to the gauge variable $\beta_{t}$ and demonstrates that (30) indeed delivers the desired extension of the PMP to stochastic optimal control problems.
Theorem 3.1.
For any choice of the gauge variable $\beta_{t}$, the resulting function $\phi_{t}(x)$ satisfies
(31) $\nabla_{x}\phi_{t}(x)=\nabla_{x}v_{t}(x)$
where $v_{t}(x)$ is the value function satisfying the HJB equation (6).
Proof.
Let us derive the evolution equation for $\nabla_{x}\phi_{t}(x)$ implied by (27):
(32a) $-\partial_{t}\nabla_{x}\phi_{t}(X_{t})$ $=D_{x}^{2}\phi_{t}(X_{t})\dot{X}_{t}-\dot{P}_{t}$
(32b) $=D_{x}^{2}\phi_{t}(X_{t})\left(b(X_{t})-G(X_{t})R^{-1}G(X_{t})^{\rm T}P_{t}-\beta_{t}\right)\,+$
(32c) $\qquad\frac{1}{2}\nabla_{x}(\Sigma:D_{x}^{2}\phi_{t}(X_{t}))+\nabla_{x}c(X_{t})-\frac{1}{2}\nabla_{x}\|RG(X_{t})^{\rm T}P_{t}\|_{R}^{2}\,+$
(32d) $\qquad\left(D_{x}b(X_{t})\right)^{\rm T}P_{t}+D_{x}^{2}\phi_{t}(X_{t})\beta_{t}$
(32e) $=D_{x}^{2}\phi_{t}(X_{t})\,b(X_{t})-\frac{1}{2}\nabla_{x}\|RG(X_{t})^{\rm T}\nabla_{x}\phi_{t}(X_{t})\|^{2}_{R}+\nabla_{x}c(X_{t})\,+$
(32f) $\qquad\frac{1}{2}\nabla_{x}(\Sigma:D_{x}^{2}\phi_{t}(X_{t}))+\left(D_{x}b(X_{t})\right)^{\rm T}\nabla_{x}\phi_{t}(X_{t}).$
Hence, since it follows that $\tilde{y}_{t}:=\nabla_{x}\phi_{t}$ satisfies the reverse time partial differential equation
(33) $-\partial_{t}\tilde{y}_{t}=\nabla_{x}(\tilde{y}_{t}\cdot b)+\frac{1}{2}\nabla_{x}(\Sigma:D_{x}\tilde{y}_{t})+\nabla_{x}c-\frac{1}{2}\nabla_{x}\|RG^{\rm T}\tilde{y}_{t}\|^{2}_{R}$
and $\tilde{y}_{T}=\nabla_{x}f$, we find from (8) that $\tilde{y}_{t}=y_{t}$ regardless of the choice of the gauge function $\beta_{t}$. ∎
In line with (15), a natural choice for the gauge function is
(34) $\beta_{t}=\frac{1}{2}\Sigma\nabla_{x}\log\rho_{t}(X_{t}),$
where $\rho_{t}(x)$ denotes the law of $X_{t}$ as defined by (13). This choice leads to
(35) $\mathcal{L}_{\beta_{t}}g=-\frac{1}{2}\Sigma\nabla_{x}\log\rho_{t}\cdot\nabla_{x}g+\frac{1}{2}\Sigma:D_{x}^{2}g,$
which is the generator of a reversible diffusion process with invariant measure $\rho_{t}$ [19]. Since $\mathcal{L}_{\beta_{t}}^{\dagger}\rho_{t}=0$ for the dual operator $\mathcal{L}_{\beta_{t}}^{\dagger}$ [19], this is also the choice of $\beta_{t}$ which makes the variation with respect to $\phi_{t}$ vanish. We recall that this choice of $\beta_{t}$ leads to $\rho_{t}=\pi_{t}$ with $\pi_{t}$ the law of $X_{t}$ arising from the SDE (1).
Alternatively, consider
(36) $\beta_{t}=-G(X_{t})RG(X_{t})^{\rm T}P_{t},$
which eliminates the control from the evolution equation in $X_{t}$ since
(37) $G(X_{t})U_{t}=-G(X_{t})RG(X_{t})^{\rm T}P_{t}=\beta_{t}.$
Hence a single forward and reverse time simulation of
(38a) $\dot{X}_{t}$ $=b(X_{t}),$
(38b) $-\dot{P}_{t}$ $=\left(D_{x}b(X_{t})\right)^{\rm T}P_{t}-\frac{1}{2}\nabla_{x}\|RG(X_{t})^{\rm T}\nabla_{x}\phi_{t}(X_{t})\|^{2}_{R}\,+$
(38c) $\qquad\qquad\nabla_{x}c(X_{t})+\frac{1}{2}\nabla_{x}(\Sigma:D_{x}^{2}\phi_{t}(X_{t}))$
subject to $X_{0}=a\sim p_{0}$ and $P_{T}(a)=\nabla_{x}f(X_{T}(a))$ solves the stochastic optimal control problem. We note that such a decoupling is possible even when $\Sigma$ is singular or even zero. This is in difference to the classical as well as the stochastic PMP [2].
Both choices for the gauge variable can also be combined into
(39) $\beta_{t}=\frac{1}{2}\Sigma\nabla_{x}\log\rho_{t}(X_{t})-G(X_{t})RG(X_{t})^{\rm T}P_{t}-G(X_{t})u_{t}^{\rm ref}(X_{t}),$
where $u_{t}^{\rm ref}(x)$ is a known reference control; if available. In the context of NMPC [23, 7], such a control could arise from the last receding horizon control window.
4. Discounted cost optimal control problems
We now consider infinite horizon optimal control problems with discounted cost function (3). Such problems can be dealt with by simply introducing the damping term $-\gamma P_{t}$ into the evolution equation for the momentum variable $P_{t}$ [2]. The boundary condition for $P_{T}$ is replaced by the transversality condition
(40) $\lim_{T\to\infty}e^{-\gamma T}P_{T}(a)\cdot X_{T}(a)=0,$
which follows from (21) adapted to the case of discounted cost; i.e.
(41) $\mathcal{S}(X,P,\beta)=\int_{\mathbb{R}^{d_{x}}}\left\{\int_{0}^{\infty}e^{-\gamma t}\left(P_{t}^{\rm T}\dot{X}_{t}-H(X_{t},P_{t},\beta_{t})\right){\rm d}t\right\}p_{0}{\rm d}a,$
and the fact that boundary terms arising from integration by parts should vanish.
A formal derivation of the action function (41) can again start from a reformulation of the associated HJB equation:
(42a) $-\partial_{t}v_{t}$ $=-\gamma v_{t}+b\cdot\nabla_{x}v_{t}+\frac{1}{2}\nabla_{x}\cdot(\Sigma\nabla_{x}\phi_{t})+c-\frac{1}{2}\|RG(x)^{\rm T}\nabla_{x}v_{t}\|^{2}_{R},$
(42b) 00 0 $=\phi_{t}-v_{t}.$
The second step is to introduce Lagrange multipliers $\rho_{t}$ and $\beta_{t}$ in order to enforce (42) in terms of the action functional
(43a) $\mathcal{S}(\tilde{v},\rho,\beta)$ $=-\int_{0}^{T}\int_{\mathbb{R}^{d_{x}}}\left\{\partial_{t}\tilde{v}_{t}+\nabla_{x}\tilde{v}_{t}\cdot b+e^{-\gamma t}c-\frac{e^{\gamma t}}{2}\|RG^{\rm T}\nabla_{x}\tilde{v}_{t}\|^{2}_{R}\right\}\rho_{t}{\rm d}x\,{\rm d}t\,+$
(43b) $\qquad\qquad\int_{0}^{T}\int_{\mathbb{R}^{d_{x}}}\left\{\beta_{t}\cdot\nabla_{x}(\tilde{\phi}_{t}-\tilde{v}_{t})+\frac{1}{2}\nabla_{x}\cdot(\Sigma\nabla_{x}\tilde{\phi}_{t})\right\}\rho_{t}{\rm d}x\,{\rm d}t.$
Here we have scaled the value function by $e^{-\gamma t}$; i.e. $\tilde{v}_{t}=e^{-\gamma t}v_{t}$. After transferring to the Lagrangian framework, we undo the scaling in $\tilde{P}_{t}=\nabla_{x}\tilde{v}_{t}$ to obtain the action functional (41).
Since we are interested in equilibrium states only; i.e. $t\to\infty$, we can run both equations forward in time resulting in
(44a) $\dot{X}_{t}$ $=\nabla_{p}H(X_{t},P_{t},\beta_{t}),$
(44b) $\dot{P}_{t}$ $=-\gamma P_{t}+\nabla_{x}H(X_{t},P_{t},\beta_{t}).$
Then $(X_{\infty}(a),P_{\infty}(a))$ define the optimal control law (28) via the implied $\nabla_{x}\phi_{\infty}(x)$. In this setting, the choice (34) for $\beta_{t}$ emerges as the most natural one.
We note that such a joint forward integration is not possible for forward-backward SDE formulations of the PMP [2].
5. Numerical implementation
Numerical implementations of (30) rely on a set of particles $(X_{t}^{(i)},P_{t}^{(i)})$, $i=1,\ldots,M$, which evolve according to
(45a) $\dot{X}_{t}^{(i)}$ $=\nabla_{P_{i}}\mathcal{H}(\{X_{t}^{(j)}\},\{P_{t}^{(j)}\}),$
(45b) $-\dot{P}_{t}^{(i)}$ $=\nabla_{X_{i}}\mathcal{H}(\{X_{t}^{(j)}\},\{P_{t}^{(j)}\})$
for appropriate Hamiltonian $\mathcal{H}$ and subject to boundary conditions
(46) $X_{0}^{(i)}=a_{i}\sim\pi_{0},\qquad P_{T}^{(i)}=-\nabla_{x}f(X_{T}^{(i)}).$
In order to get closed form ODEs, we need to specify $\mathcal{H}$ which involves approximating $\phi_{t}(x)$ as defined by the constraint (27). We discuss a few choices below. These approximations apply equally well to the discounted cost mean-field PMP formulation from Section 4.
5.1. A regression formulation
The constraint (27) can be approximated via the method of least squares. For example, one could use
(47) $\mathcal{V}_{t}(\gamma_{t})=\frac{1}{2M}\sum_{i=1}^{M}\|P_{t}^{(i)}+\nabla_{x}\tilde{\phi}(X_{t}^{(i)};\gamma_{t})\|^{2}$
for suitable class of functions $\tilde{\phi}(\cdot;\gamma)$ parametrized by $\gamma\in\mathbb{R}^{d_{\gamma}}$. This approach is not strictly variational unless $\Sigma=0$; but it can be applied to any choice of the gauge variable $\beta_{t}(a)$. See Section 5.3 below for a strictly variational approach.
As a specific example, consider
(48) $\beta_{t}^{(i)}=\frac{1}{2}G(X_{t}^{(i)})R^{-1}G(X_{t}^{(i)})(A_{t}X_{t}^{(i)}+c_{t}-P_{t}^{(i)})$
together with a bilinear ansatz
(49) $\tilde{\phi}(x;\gamma)=\frac{1}{2}x^{\rm T}Ax+x^{\rm T}c,$
which makes the term depending on $\Sigma$ in (22) vanish. Thus the Hamiltonian (22) gives rise to
(50a) $\mathcal{H}_{t}(\{X_{t}^{(j)}\},\{P_{t}^{(i)}\})$ $=\sum_{i=1}^{M}\left\{-\frac{1}{2}\|RG(X_{t}^{(i)})^{\rm T}P_{t}^{(i)}\|^{2}_{R}+(P_{t}^{(i)})^{\rm T}b(X_{t}^{(i)})+c(X_{t}^{(i)})\,+\right.$
(50b) $\qquad\qquad\left.\frac{1}{2}\|RG(X_{t}^{(i)})^{\rm T}(A_{t}X_{t}^{(i)}-c_{t}-P_{t}^{(i)})\|^{2}_{R}\right\}$
and the equations of motion (45) reduce to
(51a) $\dot{X}_{t}^{(i)}$ $=b(X_{t}^{(i)})-G(X_{t}^{(i)})RG(X_{t}^{(i)})^{\rm T}\left(A_{t}X_{t}^{(i)}+c_{t}\right),$
(51b) $-\dot{P}_{t}^{(i)}$ $=(D_{x}b(X_{t}^{(i)}))^{\rm T}P_{t}^{(i)}+\nabla_{x}c(X_{t}^{(i)})+\frac{1}{2}\nabla_{x}\|RG(X_{t}^{(i)})^{\rm T}(A_{t}X_{t}^{(i)}+c_{t})\|_{R}^{2}\,-$
(51c) $\qquad\qquad\nabla_{x}(G(X_{t}^{(i)})P_{t}^{(i)})^{\rm T}RG(X_{t}^{(i)})^{\rm T}(A_{t}X_{t}^{(i)}+c_{t})$
with $A_{t}$ and $c_{t}$ determined by
(52) $A_{t}=C_{t}^{px}(C_{t}^{xx})^{-1},\qquad c_{t}=m^{p}_{t}-A_{t}m^{x}_{t}.$
Here $C_{t}^{px}$ and $C_{t}^{xx}$ denote the empirical covariance matrices between $\{P_{t}^{(i)}\}$ and $\{X_{t}^{(i)}\}$, and $\{X_{t}^{(i)}\}$ with itself, respectively. The empirical mean of $\{X_{t}^{(i)}\}$ and $\{P_{t}^{(i)}\}$, respectively, is denoted by $m_{t}^{x}$ and $m_{t}^{p}$, respectively. Note that $A_{t}$ is not symmetric in general; but can be made symmetric via $1/2(A_{t}^{\rm T}+A_{t})$.
Formulation (51) simplifies further in case $G$ does not depend on $x$; i.e.
(53a) $\dot{X}_{t}^{(i)}$ $=b(X_{t}^{(i)})-GRG^{\rm T}\left(A_{t}X_{t}^{(i)}+c_{t}\right),$
(53b) $-\dot{P}_{t}^{(i)}$ $=(D_{x}b(X_{t}^{(i)}))^{\rm T}P_{t}^{(i)}+\nabla_{x}c(X_{t}^{(i)})+A^{\rm T}G^{\rm T}RG^{\rm T}(A_{t}X_{t}^{(i)}+c_{t}-P_{t}^{(i)}).$
While the regression approach can be combined with any choice of $\beta_{t}$, the following two approaches rely on the specific choice (34).
5.2. Discrete Schrödinger bridge approximation
Let us consider (26) with $\beta_{t}$ given by (34) in more detail. We replace the resulting generator (35) by a discrete Schrödinger bridge approximation $m_{t}^{(ij)}$ [16, 28, 8] over the set of particles $\{X_{t}^{(j)}\}$ giving rise to
(54a) ($\mathcal{H}(\{X_{t}^{(j)}\},\{P_{t}^{(j)}\})$ $=\frac{1}{M}\sum_{i=1}^{M}\left\{-\frac{1}{2}\|RG(X_{t}^{(i)})^{\rm T}P_{t}^{(i)}\|_{R}^{2}+P_{t}^{(i)}\cdot b(X_{t}^{(i)})\right\}\,+$
(54b) $\qquad\qquad\sum_{i=1}^{M}\left\{G_{t}^{(i)}(\{X_{t}^{(j)}\})+\beta_{t}^{(i)}\cdot P_{t}^{(i)}+c(X_{t}^{(i)})\right\}$
with
(55) $\beta_{t}^{(i)}:=-\sum_{i=1}^{M}m_{t}^{(ij)}X_{t}^{(j)}$
and
(56) $G_{t}^{(i)}(\{X_{t}^{(j)}\}):=\sum_{j=1}^{M}m_{t}^{(ij)}\phi_{t}(X_{t}^{(j)}).$
The coefficients $m_{t}^{(ij)}\geq 0$ are defined by
(57) $m_{t}^{(ij)}=\frac{1}{\varepsilon}\left(v_{t}^{(i)}d_{t}^{(ij)}v_{t}^{(j)}-\delta_{ij}\right),\qquad d_{t}^{(ij)}=e^{-\frac{1}{2\varepsilon}\|X_{t}^{(i)}-X_{t}^{(j)}\|_{\Sigma}^{2}},$
with $\{v_{t}^{(i)}\}$ chosen such that
(58) $\sum_{i}m_{t}^{(ij)}=\sum_{j}m_{t}^{(ij)}=0.$
Here $\delta_{ij}$ denotes the Kronecker delta function; i.e., $\delta_{ij}=1$ if $i=j$ and $\delta_{ij}=0$ otherwise. The regularization parameter $\varepsilon>0$ needs to be chosen appropriately [28]. A particularly elegant formulation arises from the choice $\varepsilon=\Delta t$ with $\Delta t>0$ the step-size of an Euler discretization of the Hamiltonian equations of motion [8]. A fast iteterative algorithm for computing the coefficients $v_{t}^{(i)}\geq 0$ has been given in [16, 28].
Writing down the associated Hamiltonian equations of motion (45), we obtain a fully variational implementation of the proposed PMP. However, computing gradients of $m_{t}^{(ij)}$ with respect to particle positions $X_{t}^{(k)}$ is non-trivial and we use the simpler approximation
(59a) $\dot{X}_{t}^{(i)}$ $=b(X_{t}^{(i)})-G(X_{t}^{(i)})RG(X_{t}^{(i)})^{\rm T}P_{t}^{(i)}-\sum_{j=1}^{M}m_{t}^{(ij}X_{t}^{(j)},$
(59b) $-\dot{P}_{t}^{(i)}$ $=(Db(X_{t}^{(i)})^{\rm T}P_{t}^{(i)}+\nabla_{x}c(X_{t}^{(i)})-\frac{1}{2}\nabla_{x}\|RG(X_{t})^{\rm T}P_{t}\|^{2}_{R}+\sum_{j=1}^{M}m_{t}^{(ij)}P_{t}^{(j)}.$
The discrete Schrödinger bridge approximation does not require an explicit approximation of $\phi_{t}(x)$, which simplifies numerical implementations. If $\nabla_{x}\phi_{t}(x)$ is needed it can be obtained via (47).
5.3. Variational approach
A particularly versatile approach is to approximate $\beta_{t}$ and $\phi_{t}$ via a parametric ansatz $\beta_{t}(x)=\tilde{\beta}(x;\gamma)$ and $\phi_{t}(x)=\tilde{\phi}(x;\alpha_{t})$, respectively, and to consider the empirical Hamiltonian
(60a) ($\mathcal{H}(\{X_{t}^{(j)}\},\{P_{t}^{(j)}\},\gamma,\alpha)\,=$
(60b) $\qquad\qquad\sum_{j=1}^{M}\left\{-\frac{1}{2}\|RG(X_{t}^{(j)})^{\rm T}P_{t}^{(j)}\|^{2}_{R}+P_{t}^{(j)}\cdot b(X_{t}^{(j)})+c(X_{t}^{(j)})\,+\right.$
(60c) $\qquad\qquad\left.\tilde{\beta}_{t}(X_{t}^{(j)},\gamma)\cdot\nabla_{x}\left(\tilde{\phi}(X_{t}^{(j)},\alpha)-X_{t}^{(j)}\cdot P_{t}^{(j)}\right)+\frac{1}{2}\Sigma:D_{x}^{2}\tilde{\phi}(X_{t}^{(j)};\alpha)\right\}$
with equations of motion
(61a) $\dot{X}_{t}^{(i)}$ $=\nabla_{X^{(i)}}\mathcal{H}(\{X_{t}^{(j)}\},\{P_{t}^{(j)}\},\gamma_{t},\alpha_{t}),$
(61b) $-\dot{P}_{t}^{(i)}$ =($=\mathcal{H}(\{X_{t}^{(j)}\},\{P_{t}^{(j)}\},\gamma_{t},\alpha_{t}),$
(61c) $(\gamma_{t},\alpha_{t})$ $=\arg\min_{(\gamma,\alpha)}\mathcal{H}(\{X_{t}^{(j)}\},\{P_{t}^{(j)}\},\gamma,\alpha).$
Appropriate choices for $\tilde{\beta}$ and $\tilde{\phi}$ can follow the approaches considered, for exmaple, in [15].
6. Numerical experiments
We present results from two numerical examples. Both cases involve infinite horizon cost functions of the form (3) and controlled ODEs. We employ the initial value formulation (44) and its numerical approximation using discrete Schrödinger bridges as outlined in Section 5.2. In both cases, we add a diffusion (viscosity) regularization in (22) in line with the concept of viscosity solutions to the associated HJB equations and the vanishing viscosity method [3].
6.1. Inverted pendulum
Figure 1. Upper right panel: Time evolution of the ensemble mean angle and ensemble mean velocity. Upper left panel: Control law (65) at final time as function of $(\theta,v)$. Lower right panel: Control law (65) at final time as function of $\theta$. Lower left panel: Control law (65) at final time as function of $v$.
We consider a controlled inverted pendulum with friction [18]. The state variable $X=(\theta,v)^{\rm T}$ is two-dimensional with equations of motion
(62a) $\dot{\theta}_{t}$ $=v_{t},$
(62b) $\dot{v}_{t}$ $=\sin(\theta_{t})-\sigma v_{t}+\cos(\theta_{t})U_{t}$
and $\sigma=5$. Consider the running cost function
(63) $c(x)=\frac{\alpha}{2}\left(\theta^{2}+v^{2}\right)$
with $\alpha=25$ and set $R=1$ in (4). We use the discount factor $\gamma=1$ in (3). Note that $(0,0)$ is an unstable equilibrium point under $U_{t}\equiv 0$.
In order to implement and, in fact, regularize the mean-field PMP, we add diffusion (viscosity) by setting $\Sigma=0.1I$ in (22). We initialize the particles at the stable equilibrium point $x_{\rm s}=(\pi,0)$ with added Gaussian noise of variance $0.1I$; i.e. $X_{0}^{(i)}\sim{\rm N}(x_{\rm s},0.1I)$. The initial momenta are set to $P_{0}^{(i)}=-2SX_{0}^{(i)}$ with
(64) $S=\left(\begin{array}[]{cc}0&1\\ 1&0\end{array}\right).$
The discrete Schrödinger bridge based formulation from Section 5.2 has been implemented with ensemble size $M=200$ and $\varepsilon=\Delta t$. The resulting evolution equations (44) have been simulated over 8000 Euler time-steps with time-step $\Delta t=0.025$.
The results can be found in Figure 1. It can be clearly seen that the mean-field formulation leaves the stable equilibrium and rapidly equilibrates at the unstable equilibrium. The resulting final control terms
[…]
strongly depend on $\theta$ while the dependence on $v$ is less structured.
We also note that keeping a non-zero diffusion matrix $\Sigma=0.1I$ in (22) leads to a desirable exploration of state space and, hence, to a stabilization/regularization of the control problem. However, a wider exploration of state space implies the need for larger ensemble sizes. Hence a balance of computational cost and state space exploration via added diffusion is required in practice and needs to be explored further.
6.2. Lorenz-63
Figure 2. Controlled Lorenz-63 model: Displayed is the ${\rm x}$-component of a single particle trajectory chosen at random from the ensemble $\{X_{t}^{(i)}\}$ and its associated control $U_{t}^{(i)}$ as a function of time.
We follow [12] and consider an optimal control problem for the Lorenz-63 model
(66) $\dot{X}_{t}=f(X_{t})+U_{t},\qquad f(X)=\left(\begin{array}[]{l}\sigma({\rm y}-{\rm x})\\ -{\rm x}{\rm z}+r{\rm x}-{\rm y}\\ {\rm x}{\rm y}-b{\rm z}\end{array}\right),$
in the state variable $X=({\rm x},{\rm y},{\rm z})^{\rm T}\in\mathbb{R}^{3}$ and with parameters $\sigma=10$, $r=28$, and $b=8/3$ [13]. The control problem is to restrict solutions of (66) to ${\rm x}_{t}\geq 0$ and the authors of [12] introduce the running cost
(67) $c(X)=\frac{\alpha}{2}\left(\min({\rm x},0)\right)^{2}.$
In our experiments, we have followed this cost function and have set $\alpha=5000$ and $R=1$ in (4). Contrary to the setting of [12], we have implemented a discounted cost function (3) with $\gamma=1$ instead of a receding horizon NMPC approach [7, 23].
We have implemented the Schrödinger bridge based mean-field PMP formulation with $\varepsilon=0.1$, step-size $\Delta t=0.001$, and $M=100$ particles. We again added artificial diffusion (viscosity) via $\Sigma=0.5I$ in (22). The results can be found in Figures 2 and 3. The impact of the control can be clearly detected. In particular, by enforcing ${\rm x}_{t}\geq 0$, the chaotic nature of the Lorenz-63 model has been eliminated. The initial spiral towards to quasi-periodic attractor is due to the chosen initialization
(68) $X_{0}=(7.8590;7.1136;27.2293)^{\rm T}+\Xi_{0},\qquad\Xi_{0}\sim{\rm N}(0,0.1I),$
of the Lorenz-63 problem. Our results are in qualitative agreement with the findings from [12] which are based on a combination of data assimilation [26, 6] and NMPC [7, 23].
Figure 3. Controlled Lorenz-63 model: Displayed is the three-dimensional trajectory of a single particle chosen at random from the controlled ensemble $\{X_{t}^{(i)}\}$.
7. Conclusions
A deterministic mean-field formulation of the PMP has been presented. Similar to the well-known forward-backward SDE formulation of the stochastic PMP [2] and its numerical implementation; see e.g. [5], a regression problem (27) arises when discretized as an interacting particle system. Several numerical implementations have been discussed. Contrary to its stochastic counterpart, the mean-field approach leads to smooth particle trajectories which should help with reducing the variance in the involved approximations. Furthermore, the freedom in the choice of the gauge variable $\beta_{t}$ allows for a decoupling of the forward and backward ODEs in $X_{t}$ and $P_{t}$, respectively. Furthermore, a forward simulation of the coupled Hamiltonian ODEs is possible in case of infinite horizon optimal control problems. Numerical results have been presented for a controlled inverted pendulum as well as a controlled Lorenz-63 system. While the numerical experiments focus on infinite horizon control problems and the Schrödinger bridge approximation, application of the proposed methodology to NMPC [23, 7] and variational approximations will be investigated in future work. Another area of interest is provided by extensions to partially observed optimal control problems [22, 1, 17, 29, 24].
Acknowledgment. This work has been funded by Deutsche Forschungsgemeinschaft (DFG) - Project-ID 318763901 - SFB1294.

### v3

@@
 Contrary to the well-known stochastic Pontryagin minimum principle involving forward-backward stochastic differential equations, the proposed formulation is deterministic and of mean-field type.
-The Hamiltonian structure of the proposed Pontryagin minimum principle is achieved via the introduction of an appropriate gauge variable.
+We denote it the McKean–Pontryagin minimum principle.
+The Hamiltonian structure of the proposed McKean–Pontryagin minimum principle is achieved via the introduction of an appropriate gauge variable.
 The gauge freedom can be used to decouple the forward and reverse time equations; hence simplifying the solution of the underlying boundary value problem.
@@
 In this case, the mean-field formulation allows converting the computation of the desired optimal control law into solving a pair of forward mean-field ordinary differential equations.
-The proposed mean-field formulation of the Pontryagin minimum principle is tested numerically for a controlled inverted pendulum and a controlled Lorenz-63 system.
+The McKean–Pontryagin minimum principle is tested numerically for a controlled inverted pendulum and a controlled Lorenz-63 system.
 1. Introduction
@@
 Here $B_{t}$ denotes $d_{x}$-dimensional Brownian motion, $\Sigma\in\mathbb{R}^{d_{x}\times d_{x}}$ the symmetric positive definite diffusion matrix, $R\in\mathbb{R}^{d_{u}\times d_{u}}$ a symmetric positive definite weight matrix, $G(x)\in\mathbb{R}^{d_{x}\times d_{u}}$ the possibly position-dependent control matrix, $c(x)\geq 0$ the running cost, $f(x)\geq 0$ the terminal cost, and $\gamma>0$ the discount factor.
-See, for example, [2] for more details.
+See, for example, [5, 24] for more details.
 We also introduce the weighted norm $\|\cdot\|_{R}$ via
 (4) $\|u\|^{2}_{R}=u^{\rm T}R^{-1}u.$
-The aim is to find the closed loop control law $U_{t}=u_{t}(X_{t})$ that minimizes the cost $J_{T}$ over the set of admissible control laws.
-It is well known, that the associated optimal control problems for ordinary differential equations (ODEs); i.e.
+The aim is to find the closed loop control law $U_{t}=u_{t}(X_{t})$ that minimizes the cost (2) or (3), respectively, over the set of admissible control laws.
+It is well known that the associated optimal control problems for ordinary differential equations (ODEs); i.e.
 $\Sigma=0$ in (1), can be solved by the Pontryagin minimum (maximum) principle (PMP) [28].
-The PMP has been extended to SDEs (1) in terms of forward-backward SDEs [20, 2].
+The PMP has been extended to SDEs (1) in terms of forward-backward SDEs [27, 5, 2].
 In this paper, we provide an alternative extension via a deterministic mean-field formulation, which leads to interacting particle implementations in the form of Hamiltonian ODEs.
 The formulation is attractive from a computational perspective since it allows for a decoupling of the forward and backward evolution equations for finite horizon optimal control problems and a forward simulation for infinite horizon optimal control problems.
-The proposed mean-field approach also delivers closed loop control laws in contrast to the classical PMP for ODEs [21].
-The proposed methodology can then also be used to approximate viscosity solutions for controlled ODEs by setting $\Sigma=\epsilon I$ for sufficiently small $\epsilon$ [3].
-Mean-field formulations and their interacting particle approximations have been considered before in the context of the Hamilton–Jacobi–Bellman (HJB) equation [2].
+Our mean-field approach delivers closed loop control laws in contrast to the classical PMP for ODEs [28].
+The proposed methodology can then also be used to approximate viscosity solutions for controlled ODEs by setting $\Sigma=\epsilon I$ for sufficiently small $\epsilon>0$ [6].
+Mean-field formulations and their interacting particle approximations have been considered before in the context of the Hamilton–Jacobi–Bellman (HJB) equation [5, 2].
 The key idea is to connect the time evolution of the value function to a probability density function evolving under an appropriate nonlinear Fokker–Planck equation and its mean-field representation [20, 16, 32, 17].
 In this paper, we directly approximate the gradient of the value function and derive appropriate Hamiltonian evolution equations; thus extending mean-field formulations to the PMP.
+We denote our mean-field formulation the McKean–Pontryagin minimum principle.
+As a further alternative, we mention the application of direct optimization methods to learn a parametric form of the desired control laws in the form of $u_{t}(x;\theta)$ for finite horizon cost functions (2).
+See, for example, [8, 9, 25, 15], which rely on deep learning approximation architectures.
+Similarly, infinite horizon discounted cost problems can also be reformulated as optimization problems in parameters $\theta$ based on approximations of the underlying stationary HJB equation; an approach pioneered in reinforcement learning [24].
 The remainder of the paper is structured as follows.
@@
 A number of numerical approximation techniques are discussed in Section 6; including an approximation based on discrete Schrödinger bridges [22, 35] as outlined in Section 6.2.
-The mean-field formulation is tested for two infinite horizon discounted cost (3) optimal control problems in Section 6.
-The first example is provided by a controlled inverted pendulum [18] while the second example is provided by a controlled Lorenz-63 system [12].
+The McKean–Pontryagin formulation is tested for two infinite horizon discounted cost (3) optimal control problems in Section 7.
+The first example is provided by a controlled inverted pendulum [24] while the second example concerns a controlled Lorenz-63 system [18].
+The final numerical example from Section 7.3 presents results from an implementation of the McKean–Pontryagin formulation in a receding horizon setting applied again to a controlled inverted pendulum.
 2. Mathematical background
@@
 in $(X_{t},P_{t})$ subject to the boundary conditions $X_{0}=a$ and $P_{T}=\nabla_{x}f(X_{T})$ [28].
-Formulation (7) offers significant computational savings compared to (6) and has been widely used especially in the context of nonlinear model predictive control (NMPC) [7, 23].
-The PMP for ODEs has been extended to SDEs (1) giving rise to the, so called, stochastic PMP with (7) replaced by a pair of forward-backward SDEs [2].
+Formulation (7) offers significant computational savings compared to (6) and has been used especially in the context of nonlinear model predictive control (NMPC) [11, 30].
+The PMP for ODEs has been extended to SDEs (1) giving rise to the, so called, stochastic PMP with (7) replaced by a pair of forward-backward SDEs [5, 2].
 In this paper, we develop a Hamiltonian mean-field formulation in the spirit of the PMP-based Hamiltonian ODE formulation (7), which, in turn, can be approximated by interacting particle systems.
@@
 In this case, adding diffusion via $\Sigma=\epsilon I$ with $\epsilon\to 0$ sufficiently small allows one to numerically approximate viscosity solutions of (6) [6].
-3. Pontryagin mean-field formulation
+3. McKean–Pontryagin Minimum Principle
 In this section, we develop a novel mean-field approach to the PMP starting from an Eulerian perspective in the spirit of ideal fluid dynamics [34].
 Specifically, we combine the HJB equation (6) and a Liouville equation for the flow of densities $\rho_{t}(x)$ under a single variational principle.
-In order to do so, we assume that the initial states $X_{0}$ are distributed according to a given distribution $p_{0}(x)$.
-Following [27], we then transform the variational principle into a Lagrangian framework in terms of evolving positions $X_{t}\in\mathbb{R}^{d_{x}}$ and momenta $P_{t}\in\mathbb{R}^{d_{x}}$.
-These equations are Hamiltonian and provide the novel PMP formulation.
+In order to do so, we assume that the initial states $X_{0}=a$ in (1) are no longer fixed but distributed instead according to a given distribution $p_{0}(x)$.
+Following [34], we then transform the variational principle into a Lagrangian framework in terms of evolving positions $X_{t}\in\mathbb{R}^{d_{x}}$ and momenta (co-states) $P_{t}\in\mathbb{R}^{d_{x}}$.
+These equations are Hamiltonian and provide the novel McKean–Pontryagin formulation.
 While we consider the finite horizon optimal control problem (2) in this section; an extension to infinite horizon optimal control problems can be found in Section 4.
@@
 (10a) $-\partial_{t}v_{t}$ $=b\cdot\nabla_{x}v_{t}+\frac{1}{2}\Sigma:D_{x}^{2}\phi_{t}+c-\frac{1}{2}\|RG(x)^{\rm T}\nabla_{x}v_{t}\|^{2}_{R},$
-(10b) 00 0 $=\phi_{t}-v_{t}.$
+(10b) 0 0 $=\phi_{t}-v_{t}.$
 The second step is to introduce Lagrange multipliers $\rho_{t}(x)$ and $\beta_{t}(x)$ in order to enforce (10) in terms of the action functional
-(11a) $\mathcal{S}(v,\rho,\beta)$ $=-\int_{0}^{T}\int_{\mathbb{R}^{d_{x}}}\left\{\partial_{t}v_{t}+\nabla_{x}v_{t}\cdot b+c-\frac{1}{2}\|RG^{\rm T}\nabla_{x}v_{t}\|^{2}_{R}\right\}\rho_{t}{\rm d}x\,{\rm d}t\,+$
+(11a) $\mathcal{S}(v,\rho,\beta)$ $=-\int_{0}^{T}\int_{\mathbb{R}^{d_{x}}}\left\{\partial_{t}v_{t}+\nabla_{x}v_{t}\cdot b+c-\frac{1}{2}\|RG^{\rm T}\nabla_{x}v_{t}\|^{2}_{R}\right\}\rho_{t}{\rm d}x\,{\rm d}t\,-$
 (11b) $\qquad\qquad\int_{0}^{T}\int_{\mathbb{R}^{d_{x}}}\left\{\beta_{t}\cdot\nabla_{x}(\phi_{t}-v_{t})+\frac{1}{2}\nabla_{x}\cdot(\Sigma\nabla_{x}\phi_{t})\,\right\}\rho_{t}{\rm d}x\,{\rm d}t\,+$
@@
 3.2. Lagrangian formulation
-Let us now convert (11) into a corresponding Lagrangian formulation [27].
+Let us convert (11) into a corresponding Lagrangian formulation [34].
 Let $X_{\tau}(a)$ denotes the solution of (13) at time $\tau=t$ with initial condition $X_{0}=a$.
@@
 (17c) $=\int_{0}^{T}\int_{\mathbb{R}^{d_{x}}}P_{\tau}(a)\cdot\dot{X}_{\tau}(a)\,p_{0}(a){\rm d}a\,{\rm d}\tau-\int_{\mathbb{R}^{d_{x}}}\left(v_{T}(X_{T}(a))-v_{0}(X_{0}(a))\right)\,p_{0}(a){\rm d}a,$
-where we have introduced the momentum variable
+where we have introduced the momentum variable (co-state)
 (18) $P_{\tau}(a)=\nabla_{x}v_{\tau}(X_{\tau}(a)),$
@@
 (19) $\partial_{t}v_{t}(x=X_{\tau=t}(a))=\partial_{\tau}v_{\tau}(X_{\tau}(a))-\nabla_{x}v_{\tau}(X_{\tau}(a))\cdot\dot{X}_{\tau}(a)$
-as well as the fact that the total time derivative $\partial_{\tau}v_{\tau}(X_{\tau}(a))$ contributes boundary terms only. Furthermore,
+as well as the fact that the total time derivative $\partial_{\tau}v_{\tau}(X_{\tau}(a))$ contributes boundary terms only to action functional (17). Furthermore,
 (20a) $\int_{\mathbb{R}^{d_{x}}}\beta_{t}(x)\cdot\nabla_{x}(\phi_{t}(x)-v_{t}(x))\,\rho_{t}(x){\rm d}x=$
@@
 with $\tilde{\beta}_{\tau}(a)=\beta_{t}(X_{\tau}(a))$.
-The equations of motion in $(X_{\tau},P_{\tau})$ arise now from the phase space action principle [4, 9] applied to
+The equations of motion in $(X_{\tau},P_{\tau})$ arise now from the Hamiltonian phase space action principle [7, 14] applied to
 (21) $\mathcal{S}(X,P,\beta)=\int_{\mathbb{R}^{d_{x}}}\left\{\int_{0}^{T}\left(P_{\tau}\cdot\dot{X}_{\tau}-H(X_{\tau},P_{\tau},\beta_{\tau})\right){\rm d}\tau-f(X_{T})\right\}p_{0}{\rm d}a$
@@
 (22a) $H(x,p,\beta)(a)$ $:=-\frac{1}{2}\|RG(x(a))^{\rm T}p(a)\|^{2}_{R}+p(a)^{\rm T}b(x(a))+c(x(a))\,+$
-(22b) $\qquad\qquad\beta(a)\cdot\nabla_{x}\left(\phi(x(a))-x(a)^{\rm T}p(a)\right)+\frac{1}{2}\Sigma:D_{x}^{2}\phi(x(a)).$
+(22b) $\qquad\qquad\beta(a)^{\rm T}\nabla_{x}\left(\phi(x(a))-x(a)^{\rm T}p(a)\right)+\frac{1}{2}\Sigma:D_{x}^{2}\phi(x(a)).$
 The desired optimal closed loop control law (5) is provided by
@@
 The boundary conditions are $X_{0}(a)=a\sim p_{0}$ and $P_{T}(a)=\nabla_{x}f(X_{T}(a))$.
-The equations (29) are of mean-field type through $\phi_{t}(x)$, which is determined by the constraint (27).
+The equations (29) are of mean-field type through $\phi_{\tau}(x)$, which is determined by the constraint (27).
 Dropping labels $a\in\mathbb{R}^{d_{x}}$ from now on and denoting time by $t$ again, the Hamiltonian equations of motion (29) become
@@
 (30c) $-\dot{P}_{t}$ $=\left(D_{x}b(X_{t})\right)^{\rm T}P_{t}\,+$
-(30d) $\qquad\qquad\nabla_{x}\left\{c(X_{t}^{(i)})-\frac{1}{2}\|RG(X_{t})^{\rm T}P_{t}\|^{2}_{R}+\mathcal{L}_{\beta_{t}}(\phi_{t}(X_{t})-X_{t}\cdot P_{t})\right\}$
-(30e) $=\left(D_{x}b(X_{t})\right)^{\rm T}P_{t}-\frac{1}{2}\nabla_{x}\|RG(X_{t})^{\rm T}P_{t}\|^{2}_{R}+\nabla_{x}c(X_{t}^{(i)})+\mathcal{L}_{\beta_{t}}\nabla_{x}\phi_{t}(X_{t}),$
+(30d) $\qquad\qquad\nabla_{x}\left\{c(X_{t})-\frac{1}{2}\|RG(X_{t})^{\rm T}P_{t}\|^{2}_{R}+\mathcal{L}_{\beta_{t}}(\phi_{t}(X_{t})-X_{t}\cdot P_{t})\right\}$
+(30e) $=\left(D_{x}b(X_{t})\right)^{\rm T}P_{t}-\frac{1}{2}\nabla_{x}\|RG(X_{t})^{\rm T}P_{t}\|^{2}_{R}+\nabla_{x}c(X_{t})+\mathcal{L}_{\beta_{t}}\nabla_{x}\phi_{t}(X_{t}),$
 which are to be compared to the classical PMP formulation (7). The only new terms arise in connection with the generator $\mathcal{L}_{\beta_{t}}$.
@@
 For any choice of the gauge variable $\beta_{t}$, the resulting function $\phi_{t}(x)$ satisfies
-(31) $\nabla_{x}\phi_{t}(x)=\nabla_{x}v_{t}(x)$
+(31) $\nabla_{x}\phi_{t}(x)=\nabla_{x}v_{t}(x),$
 where $v_{t}(x)$ is the value function satisfying the HJB equation (6).
@@
 Since $\mathcal{L}_{\beta_{t}}^{\dagger}\rho_{t}=0$ for the dual operator $\mathcal{L}_{\beta_{t}}^{\dagger}$ [26], this is also the choice of $\beta_{t}$ which makes the variation with respect to $\phi_{t}$ vanish.
-We recall that this choice of $\beta_{t}$ leads to $\rho_{t}=\pi_{t}$ with $\pi_{t}$ the law of $X_{t}$ arising from the SDE (1).
+We recall that (15) leads to $\rho_{t}=\pi_{t}$ with $\pi_{t}$ the law of $X_{t}$ arising from the SDE (1).
 Alternatively, consider
@@
 subject to $X_{0}=a\sim p_{0}$ and $P_{T}(a)=\nabla_{x}f(X_{T}(a))$ solves the stochastic optimal control problem.
-We note that such a decoupling is possible even when $\Sigma$ is singular or even zero.
+We note that such a decoupling is possible even when $\Sigma$ is singular or indeed zero.
 This is in difference to the classical as well as the stochastic PMP [5].
@@
 (39) $\beta_{t}=\frac{1}{2}\Sigma\nabla_{x}\log\rho_{t}(X_{t})-G(X_{t})RG(X_{t})^{\rm T}P_{t}-G(X_{t})u_{t}^{\rm ref}(X_{t}),$
-where $u_{t}^{\rm ref}(x)$ is a known reference control; if available. In the context of NMPC [23, 7], such a control could arise from the last receding horizon control window.
+where $u_{t}^{\rm ref}(x)$ is a known reference control; if available. In the context of NMPC [30, 11], such a control could arise from the previous receding horizon control window.
 4. Discounted cost optimal control problems
@@
 (42a) $-\partial_{t}v_{t}$ $=-\gamma v_{t}+b\cdot\nabla_{x}v_{t}+\frac{1}{2}\nabla_{x}\cdot(\Sigma\nabla_{x}\phi_{t})+c-\frac{1}{2}\|RG(x)^{\rm T}\nabla_{x}v_{t}\|^{2}_{R},$
-(42b) 00 0 $=\phi_{t}-v_{t}.$
+(42b) 0 0 $=\phi_{t}-v_{t}.$
 The second step is to introduce Lagrange multipliers $\rho_{t}$ and $\beta_{t}$ in order to enforce (42) in terms of the action functional
@@
 $\tilde{v}_{t}=e^{-\gamma t}v_{t}$.
-After transferring to the Lagrangian framework, we undo the scaling in $\tilde{P}_{t}=\nabla_{x}\tilde{v}_{t}$ to obtain the action functional (41).
-Since we are interested in equilibrium states only; i.e. $t\to\infty$, we can run both equations forward in time resulting in
+After switching to the Lagrangian framework, we undo the scaling in $\tilde{P}_{t}=\nabla_{x}\tilde{v}_{t}$ to obtain the action functional (41).
+The associated equations of motion are
 (44a) $\dot{X}_{t}$ $=\nabla_{p}H(X_{t},P_{t},\beta_{t}),$
-(44b) $\dot{P}_{t}$ $=-\gamma P_{t}+\nabla_{x}H(X_{t},P_{t},\beta_{t}).$
-Then $(X_{\infty}(a),P_{\infty}(a))$ define the optimal control law (28) via the implied $\nabla_{x}\phi_{\infty}(x)$. In this setting, the choice (34) for $\beta_{t}$ emerges as the most natural one.
-We note that such a joint forward integration is not possible for forward-backward SDE formulations of the PMP [2].
+(44b) $-\dot{P}_{t}$ $=-\gamma P_{t}+\nabla_{x}H(X_{t},P_{t},\beta_{t}).$
+The evolution equations (44) are still posed as a boundary value problem with (44a) solved forward in time and (44b) backward in time, respectively.
+We call a solution of (44) with $\partial_{t}\phi_{t}=0$ a steady state solution and denote it by $\hat{\phi}$. Following (42), $\hat{\phi}$ satisfies the stationary HJB equation
+(45) $0=-\gamma\hat{\phi}+b\cdot\nabla_{x}\hat{\phi}+\frac{1}{2}\nabla_{x}\cdot(\Sigma\nabla_{x}\hat{\phi})+c-\frac{1}{2}\|RG(x)^{\rm T}\nabla_{x}\hat{\phi}\|^{2}_{R}.$
+Note that the positions $X_{t}$ can still undergo a time evolution; but need to satisfy
+(46) $\dot{P}_{t}=D_{x}^{2}\hat{\phi}(X_{t})\dot{X}_{t},$
+which follows from (27).
+If we are interested in approximating such steady state solutions, we can run both equations forward in time by replacing $\dot{P}_{t}$ by $-\dot{P}_{t}$ in (44) and adding an appropriate advection term, which results in
+(47a) $\dot{X}_{t}$ $=\nabla_{p}H(X_{t},\Pi_{t},\beta_{t}),$
+(47b) $\dot{\Pi}_{t}$ $=-\gamma\Pi_{t}+\nabla_{x}H(X_{t},\Pi_{t},\beta_{t})+2D_{x}^{2}\phi_{t}(X_{t})\dot{X}_{t}.$
+Here we have introduced new momenta (co-states) $\Pi_{t}$ for clarity of presentation.
+Theorem 4.1.
+Let us assume that $(X_{t},P_{t})$, $t\in\mathbb{R}$ represent a steady state solution of (44); i.e., (46) holds for suitable $\hat{\phi}$.
+Then solutions $(X_{t},\Pi_{t})$ of (47), $t\geq 0$, with initial condition $(X_{0},\Pi_{0})=(X_{0},P_{0})$ also represent a steady state solution.
+Proof.
+We recall the definition $\tilde{y}_{t}(x)=\nabla_{t}\phi_{t}(x)$ and follow the proof of Theorem 3.1. We find that (47) implies
+(48) $\dot{\Pi}_{t}=-\partial_{t}\tilde{y}_{t}(X_{t})+D_{x}^{2}\phi_{t}(X_{t})\dot{X}_{t}$
+with $\partial_{t}\tilde{y}_{t}$ now satisfying
+(49) $\partial_{t}\tilde{y}_{t}=-\gamma\tilde{y}_{t}+\nabla_{x}(\tilde{y}_{t}\cdot b)+\frac{1}{2}\nabla_{x}(\Sigma:D_{x}\tilde{y}_{t})+\nabla_{x}c-\frac{1}{2}\nabla_{x}\|RG^{\rm T}\tilde{y}_{t}\|^{2}_{R}$
+instead of (33). It follows from $\phi_{t}=\hat{\phi}$ that
+(50) $\dot{\Pi}_{t}=D_{x}^{2}\hat{\phi}(X_{t})\dot{X}_{t}$
+in agreement with (46). In other words, $(X_{t},P_{t})=(X_{t},\Pi_{t})$ for all $t\geq 0$ when initialized in steady state and steady state solutions are preserved under (47). ∎
+We close this section by mentioning that (47) is related to iterative approaches for approximating the stationary value function (or $Q$-function) used in reinforcement learning and which are based on the stationary Bellman equation [24].
+Our approach targets instead the gradients of the value function (momenta/co-states) directly.
+This has been found beneficial from a theoretical perspective [5] and provides flexibility in terms of constraints added on the states and controls.
+It is also often found that approximating the gradient directly instead of its underlying function is more robust computationally.
+Compare, for example, density versus score estimation [21, 13].
+5. Linear control problems
+We analyze the proposed formulations in the setting of controlled linear SDEs
+(51) ${\rm d}X_{t}=AX_{t}+GU_{t}+\Sigma^{1/2}{\rm d}B_{t},\qquad X_{0}\sim{\rm N}(0,C_{0})$
+for infinite time discounted cost functional (3) with running cost
+(52) $c(x)=\frac{1}{2}x^{\rm T}Cx$
+and $R=I$. We write the function $\phi_{t}(x)$ as
+(53) $\phi_{t}(x)=\frac{1}{2}x^{\rm T}\Omega_{t}x.$
+With this ansatz the Hamiltonian equations of motion (29) take the form
+(54a) $\dot{X}_{t}$ $=AX_{t}-GG^{\rm T}P_{t}-\beta_{t},$
+(54b) $-\dot{P}_{t}$ $=-\gamma P_{t}+A^{\rm T}P_{t}+CX_{t}+\Omega_{t}\beta_{t}.$
+The choice (34) for the gauge variable $\beta_{t}$ leads to
+(55) $\beta_{t}=-\frac{1}{2}\Sigma C_{t}^{-1}X_{t}$
+with $C_{t}$ the covariance matrix of the Gaussian distributed $X_{t}$ with mean zero. The constraint (27) becomes
+(56) $P_{t}=\Omega_{t}X_{t},$
+which yields
+(57a) $-\dot{\Omega}_{t}X_{t}$ $=\Omega_{t}\dot{X}_{t}-\dot{P}_{t}$
+(57b) $=\Omega_{t}\left(AX_{t}-GG^{\rm T}P_{t}-\beta_{t}\right)-\gamma P_{t}+A^{\rm T}P_{t}+CX_{t}+\Omega_{t}\beta_{t}$
+(57c) $=\left(-\gamma\Omega_{t}+\Omega_{t}A-\Omega_{t}GG^{\rm T}\Omega_{t}+A^{\rm T}\Omega_{t}+C\right)X_{t}$
+and from which we can read off the backward Riccati equation for $\Omega_{t}$ [5]:
+(58) $-\dot{\Omega}=-\gamma\Omega_{t}+\Omega_{t}A+A^{\rm T}\Omega_{t}-\Omega_{t}GG^{\rm T}\Omega_{t}+C.$
+Let us now investigate the forward evolution equations (47). We obtain
+(59a) $\dot{X}_{t}$ $=AX_{t}-GG^{\rm T}P_{t}-\beta_{t}$
+(59b) $\dot{\Pi}_{t}$ $=-\gamma\Pi_{t}+A^{\rm T}\Pi_{t}+CX_{t}+\Omega_{t}\beta_{t}+2\Omega_{t}\dot{X}_{t}$
+subject to the constraint $\Pi_{t}=\Omega_{t}X_{t}$. This time we obtain
+(60a) $\dot{\Omega}_{t}X_{t}$ $=-\Omega_{t}\dot{X}_{t}+\dot{\Pi}_{t}$
+(60b) $=\Omega_{t}\left(AX_{t}-GG^{\rm T}\Pi_{t}-\beta_{t}\right)-\gamma\Pi_{t}+A^{\rm T}\Pi_{t}+CX_{t}+\Omega_{t}\beta_{t}$
+(60c) $=\left(-\gamma\Omega_{t}+\Omega_{t}A+A^{\rm T}\Omega_{t}-\Omega_{t}GG^{\rm T}\Omega_{t}+C\right)X_{t}$
+and read off the forward Riccati equation
+(61) $\dot{\Omega}=-\gamma\Omega_{t}+\Omega_{t}A+A^{\rm T}\Omega_{t}-\Omega_{t}GG^{\rm T}\Omega_{t}+C.$
+Clearly, the forward Riccati equation is stable if and only if the backward Riccati equation is stable and their stationary solutions agree.
+Of course, our approach is to propagate $X_{t}$ and $\Pi_{t}$ and to then find the appropriate $\Omega_{t}$ by enforcing the constraint $\Pi_{t}=\Omega_{t}X_{t}$ via regression; i.e.
+(62) $\Omega_{t}=\arg\min_{\Omega\in\mathbb{R}^{d_{x}\times d_{x}}}\mathbb{E}\left[\|\Pi_{t}-\Omega X_{t}\|^{2}\right].$
+We describe numerical implementations for the general nonlinear setting in the subsequent section.
 6. Numerical implementation
-Numerical implementations of (30) rely on a set of particles $(X_{t}^{(i)},P_{t}^{(i)})$, $i=1,\ldots,M$, which evolve according to
-(45a) $\dot{X}_{t}^{(i)}$ $=\nabla_{P_{i}}\mathcal{H}(\{X_{t}^{(j)}\},\{P_{t}^{(j)}\}),$
-(45b) $-\dot{P}_{t}^{(i)}$ $=\nabla_{X_{i}}\mathcal{H}(\{X_{t}^{(j)}\},\{P_{t}^{(j)}\})$
-for appropriate Hamiltonian $\mathcal{H}$ and subject to boundary conditions
-(46) $X_{0}^{(i)}=a_{i}\sim\pi_{0},\qquad P_{T}^{(i)}=-\nabla_{x}f(X_{T}^{(i)}).$
-In order to get closed form ODEs, we need to specify $\mathcal{H}$ which involves approximating $\phi_{t}(x)$ as defined by the constraint (27).
-We discuss a few choices below.
-These approximations apply equally well to the discounted cost mean-field PMP formulation from Section 4.
+Numerical implementations of the McKean–Pontryagin formulations rely on a set of particles $(X_{t}^{(i)},P_{t}^{(i)})$, $i=1,\ldots,M$, with corresponding evolution equations which approximate the mean-field ODE formulation (30).
+At this point two principal choices arise; either one discretizes directly at the level of the mean-field ODE or, alternatively, one discretizes at the level of the Hamiltonian (22) first, giving rise to a Hamiltonian ($\mathcal{H}(\{X_{t}^{(j)}\},\{P_{t}^{(j)}\})$ with associated Hamiltonian equations of motion
+(63a) $\dot{X}_{t}^{(i)}$ $=\nabla_{P^{(i)}}\mathcal{H}(\{X_{t}^{(j)}\},\{P_{t}^{(j)}\}),$
+(63b) $-\dot{P}_{t}^{(i)}$ $=\nabla_{X^{(i)}}\mathcal{H}(\{X_{t}^{(j)}\},\{P_{t}^{(j)}\})$
+subject to boundary conditions
+(64) $X_{0}^{(i)}=a_{i}\sim\pi_{0},\qquad P_{T}^{(i)}=\nabla_{x}f(X_{T}^{(i)}).$
+Below we discuss three different particle-based approximations which primarily differ in the way the constraint (27) is dealt with.
+These approximations apply equally well to the discounted cost formulation (47) of the McKean–Pontryagin minimum principle from Section 4.
 6.1. A regression formulation
-The constraint (27) can be approximated via the method of least squares. For example, one could use
-(47) $\mathcal{V}_{t}(\gamma_{t})=\frac{1}{2M}\sum_{i=1}^{M}\|P_{t}^{(i)}+\nabla_{x}\tilde{\phi}(X_{t}^{(i)};\gamma_{t})\|^{2}$
-for suitable class of functions $\tilde{\phi}(\cdot;\gamma)$ parametrized by $\gamma\in\mathbb{R}^{d_{\gamma}}$.
-This approach is not strictly variational unless $\Sigma=0$; but it can be applied to any choice of the gauge variable $\beta_{t}(a)$.
-See Section 5.3 below for a strictly variational approach.
-As a specific example, consider
-(48) $\beta_{t}^{(i)}=\frac{1}{2}G(X_{t}^{(i)})R^{-1}G(X_{t}^{(i)})(A_{t}X_{t}^{(i)}+c_{t}-P_{t}^{(i)})$
-together with a bilinear ansatz
-(49) $\tilde{\phi}(x;\gamma)=\frac{1}{2}x^{\rm T}Ax+x^{\rm T}c,$
-which makes the term depending on $\Sigma$ in (22) vanish. Thus the Hamiltonian (22) gives rise to
-(50a) $\mathcal{H}_{t}(\{X_{t}^{(j)}\},\{P_{t}^{(i)}\})$ $=\sum_{i=1}^{M}\left\{-\frac{1}{2}\|RG(X_{t}^{(i)})^{\rm T}P_{t}^{(i)}\|^{2}_{R}+(P_{t}^{(i)})^{\rm T}b(X_{t}^{(i)})+c(X_{t}^{(i)})\,+\right.$
-(50b) $\qquad\qquad\left.\frac{1}{2}\|RG(X_{t}^{(i)})^{\rm T}(A_{t}X_{t}^{(i)}-c_{t}-P_{t}^{(i)})\|^{2}_{R}\right\}$
-and the equations of motion (45) reduce to
-(51a) $\dot{X}_{t}^{(i)}$ $=b(X_{t}^{(i)})-G(X_{t}^{(i)})RG(X_{t}^{(i)})^{\rm T}\left(A_{t}X_{t}^{(i)}+c_{t}\right),$
-(51b) $-\dot{P}_{t}^{(i)}$ $=(D_{x}b(X_{t}^{(i)}))^{\rm T}P_{t}^{(i)}+\nabla_{x}c(X_{t}^{(i)})+\frac{1}{2}\nabla_{x}\|RG(X_{t}^{(i)})^{\rm T}(A_{t}X_{t}^{(i)}+c_{t})\|_{R}^{2}\,-$
-(51c) $\qquad\qquad\nabla_{x}(G(X_{t}^{(i)})P_{t}^{(i)})^{\rm T}RG(X_{t}^{(i)})^{\rm T}(A_{t}X_{t}^{(i)}+c_{t})$
-with $A_{t}$ and $c_{t}$ determined by
-(52) $A_{t}=C_{t}^{px}(C_{t}^{xx})^{-1},\qquad c_{t}=m^{p}_{t}-A_{t}m^{x}_{t}.$
-Here $C_{t}^{px}$ and $C_{t}^{xx}$ denote the empirical covariance matrices between $\{P_{t}^{(i)}\}$ and $\{X_{t}^{(i)}\}$, and $\{X_{t}^{(i)}\}$ with itself, respectively.
-The empirical mean of $\{X_{t}^{(i)}\}$ and $\{P_{t}^{(i)}\}$, respectively, is denoted by $m_{t}^{x}$ and $m_{t}^{p}$, respectively.
-Note that $A_{t}$ is not symmetric in general; but can be made symmetric via $1/2(A_{t}^{\rm T}+A_{t})$.
-Formulation (51) simplifies further in case $G$ does not depend on $x$; i.e.
-(53a) $\dot{X}_{t}^{(i)}$ $=b(X_{t}^{(i)})-GRG^{\rm T}\left(A_{t}X_{t}^{(i)}+c_{t}\right),$
-(53b) $-\dot{P}_{t}^{(i)}$ $=(D_{x}b(X_{t}^{(i)}))^{\rm T}P_{t}^{(i)}+\nabla_{x}c(X_{t}^{(i)})+A^{\rm T}G^{\rm T}RG^{\rm T}(A_{t}X_{t}^{(i)}+c_{t}-P_{t}^{(i)}).$
+The constraint (27) can be approximated via the method of least squares. For example, one introduces the cost function
+(65) $\mathcal{V}_{t}(\eta)=\frac{1}{2M}\sum_{i=1}^{M}\|P_{t}^{(i)}-\tilde{\Phi}(X_{t}^{(i)};\eta)\|^{2}$
+for suitable class of functions $\tilde{\Phi}(\cdot;\eta)$ parametrized by $\eta\in\mathbb{R}^{d_{\eta}}$.
+The approximation $\tilde{\Phi}(\cdot;\eta_{t})$ with $\eta_{t}=\arg\min\mathcal{V}_{t}(\eta)$ is then used in (30) in place of $\nabla_{x}\phi_{t}(x)$ and gives rise to associated equations of motion in terms of the particles $(X_{t}^{(i)},P_{t}^{(i)})$, $i=1,\ldots,M$.
+Since we have directly approximated the evolution equations (30), the numerical approximation is not of the form (63) for suitable Hamiltonian $\mathcal{H}$.
+See Section 6.3 below for a strictly variational approach; i.e., where one first derives a particle approximation $\mathcal{H}$ of the Hamiltonian (22) and subsequently (63).
 While the regression approach can be combined with any choice of $\beta_{t}$, the following two approaches rely on the specific choice (34).
@@
 (66a) ($\mathcal{H}(\{X_{t}^{(j)}\},\{P_{t}^{(j)}\})$ $=\frac{1}{M}\sum_{i=1}^{M}\left\{-\frac{1}{2}\|RG(X_{t}^{(i)})^{\rm T}P_{t}^{(i)}\|_{R}^{2}+P_{t}^{(i)}\cdot b(X_{t}^{(i)})\right\}\,+$
-(54b) $\qquad\qquad\sum_{i=1}^{M}\left\{G_{t}^{(i)}(\{X_{t}^{(j)}\})+\beta_{t}^{(i)}\cdot P_{t}^{(i)}+c(X_{t}^{(i)})\right\}$
+(66b) $\qquad\qquad\frac{1}{M}\sum_{i=1}^{M}\left\{S_{t}^{(i)}(\{X_{t}^{(j)}\})+\beta_{t}^{(i)}\cdot P_{t}^{(i)}+c(X_{t}^{(i)})\right\}$
 with
@@
 and
-(56) $G_{t}^{(i)}(\{X_{t}^{(j)}\}):=\sum_{j=1}^{M}m_{t}^{(ij)}\phi_{t}(X_{t}^{(j)}).$
-The coefficients $m_{t}^{(ij)}\geq 0$ are defined by
+(68) $S_{t}^{(i)}(\{X_{t}^{(j)}\}):=\sum_{j=1}^{M}m_{t}^{(ij)}\phi_{t}(X_{t}^{(j)}).$
+The coefficients $m_{t}^{(ij)}$ are defined by
 (69) $m_{t}^{(ij)}=\frac{1}{\varepsilon}\left(v_{t}^{(i)}d_{t}^{(ij)}v_{t}^{(j)}-\delta_{ij}\right),\qquad d_{t}^{(ij)}=e^{-\frac{1}{2\varepsilon}\|X_{t}^{(i)}-X_{t}^{(j)}\|_{\Sigma}^{2}},$
-with $\{v_{t}^{(i)}\}$ chosen such that
+with $\{v_{t}^{(i)}\geq 0\}$ chosen such that
 (70) $\sum_{i}m_{t}^{(ij)}=\sum_{j}m_{t}^{(ij)}=0.$
@@
 A particularly elegant formulation arises from the choice $\varepsilon=\Delta t$ with $\Delta t>0$ the step-size of an Euler discretization of the Hamiltonian equations of motion [12].
-A fast iteterative algorithm for computing the coefficients $v_{t}^{(i)}\geq 0$ has been given in [16, 28].
-Writing down the associated Hamiltonian equations of motion (45), we obtain a fully variational implementation of the proposed PMP.
-However, computing gradients of $m_{t}^{(ij)}$ with respect to particle positions $X_{t}^{(k)}$ is non-trivial and we use the simpler approximation
-(59a) $\dot{X}_{t}^{(i)}$ $=b(X_{t}^{(i)})-G(X_{t}^{(i)})RG(X_{t}^{(i)})^{\rm T}P_{t}^{(i)}-\sum_{j=1}^{M}m_{t}^{(ij}X_{t}^{(j)},$
-(59b) $-\dot{P}_{t}^{(i)}$ $=(Db(X_{t}^{(i)})^{\rm T}P_{t}^{(i)}+\nabla_{x}c(X_{t}^{(i)})-\frac{1}{2}\nabla_{x}\|RG(X_{t})^{\rm T}P_{t}\|^{2}_{R}+\sum_{j=1}^{M}m_{t}^{(ij)}P_{t}^{(j)}.$
+A fast iterative algorithm for computing the coefficients $v_{t}^{(i)}$ has been given in [22, 35].
+Writing down the associated Hamiltonian equations of motion (63), we obtain a fully variational implementation of the proposed McKean–Pontryagin minimum principle.
+However, computing gradients of $m_{t}^{(ij)}$ with respect to particle positions $X_{t}^{(k)}$ is non-trivial and we use the simpler approximation based on the following direct application of the Schrödinger bridge approximation to (30):
+(71a) $\dot{X}_{t}^{(i)}$ $=b(X_{t}^{(i)})-G(X_{t}^{(i)})RG(X_{t}^{(i)})^{\rm T}P_{t}^{(i)}-\sum_{j=1}^{M}m_{t}^{(ij)}X_{t}^{(j)},$
+(71b) $-\dot{P}_{t}^{(i)}$ $=(Db(X_{t}^{(i)}))^{\rm T}P_{t}^{(i)}+\nabla_{x}c(X_{t}^{(i)})-\frac{1}{2}\nabla_{x}\|RG(X_{t}^{(i)})^{\rm T}P_{t}^{(i)}\|^{2}_{R}+\sum_{j=1}^{M}m_{t}^{(ij)}P_{t}^{(j)}.$
 The discrete Schrödinger bridge approximation does not require an explicit approximation of $\phi_{t}(x)$, which simplifies numerical implementations.
-If $\nabla_{x}\phi_{t}(x)$ is needed it can be obtained via (47).
+We use this approach for the numerical experiments in Sections 7.1 and 7.2.
 6.3. Variational approach
-A particularly versatile approach is to approximate $\beta_{t}$ and $\phi_{t}$ via a parametric ansatz $\beta_{t}(x)=\tilde{\beta}(x;\gamma)$ and $\phi_{t}(x)=\tilde{\phi}(x;\alpha_{t})$, respectively, and to consider the empirical Hamiltonian
-(60a) ($\mathcal{H}(\{X_{t}^{(j)}\},\{P_{t}^{(j)}\},\gamma,\alpha)\,=$
+A particularly versatile approach is to approximate $\beta_{t}$ and $\phi_{t}$ via a parametric ansatz $\beta_{t}(x)=\tilde{\beta}(x;\eta_{t})$ and $\nabla_{x}\phi_{t}(x)=\tilde{\Phi}(x;\alpha_{t})$, respectively, and to consider the Hamiltonian
+(72a) ($\mathcal{H}(\{X_{t}^{(j)}\},\{P_{t}^{(j)}\},\eta_{t},\alpha_{t})\,=$
 (72b) $\qquad\qquad\sum_{j=1}^{M}\left\{-\frac{1}{2}\|RG(X_{t}^{(j)})^{\rm T}P_{t}^{(j)}\|^{2}_{R}+P_{t}^{(j)}\cdot b(X_{t}^{(j)})+c(X_{t}^{(j)})\,+\right.$
-(60c) $\qquad\qquad\left.\tilde{\beta}_{t}(X_{t}^{(j)},\gamma)\cdot\nabla_{x}\left(\tilde{\phi}(X_{t}^{(j)},\alpha)-X_{t}^{(j)}\cdot P_{t}^{(j)}\right)+\frac{1}{2}\Sigma:D_{x}^{2}\tilde{\phi}(X_{t}^{(j)};\alpha)\right\}$
+(72c) $\qquad\qquad\left.\tilde{\beta}(X_{t}^{(j)},\eta_{t})\cdot\left(\tilde{\Phi}(X_{t}^{(j)},\alpha_{t})-P_{t}^{(j)}\right)+\frac{1}{2}\nabla_{x}\cdot\left(\Sigma\tilde{\Phi}(X_{t}^{(j)};\alpha_{t})\right)\right\}$
 with equations of motion
-(61a) $\dot{X}_{t}^{(i)}$ $=\nabla_{X^{(i)}}\mathcal{H}(\{X_{t}^{(j)}\},\{P_{t}^{(j)}\},\gamma_{t},\alpha_{t}),$
-(61b) $-\dot{P}_{t}^{(i)}$ =($=\mathcal{H}(\{X_{t}^{(j)}\},\{P_{t}^{(j)}\},\gamma_{t},\alpha_{t}),$
-(61c) $(\gamma_{t},\alpha_{t})$ $=\arg\min_{(\gamma,\alpha)}\mathcal{H}(\{X_{t}^{(j)}\},\{P_{t}^{(j)}\},\gamma,\alpha).$
-Appropriate choices for $\tilde{\beta}$ and $\tilde{\phi}$ can follow the approaches considered, for exmaple, in [15].
+(73a) $\dot{X}_{t}^{(i)}$ $=\nabla_{P^{(i)}}\mathcal{H}(\{X_{t}^{(j)}\},\{P_{t}^{(j)}\},\eta_{t},\alpha_{t}),$
+(73b) $-\dot{P}_{t}^{(i)}$ $=\nabla_{X^{(i)}}\mathcal{H}(\{X_{t}^{(j)}\},\{P_{t}^{(j)}\},\eta_{t},\alpha_{t}),$
+(73c) 0 0 $=\nabla_{\eta}\mathcal{H}(\{X_{t}^{(j)}\},\{P_{t}^{(j)}\},\eta_{t},\alpha_{t})$
+(73d) 0 0 $=\nabla_{\alpha}\mathcal{H}(\{X_{t}^{(j)}\},\{P_{t}^{(j)}\},\eta_{t},\alpha_{t}).$
+Appropriate choices for $\tilde{\beta}$ and $\tilde{\Phi}$ can follow the approaches considered, for example, in [21]. Below, we discuss the case of linear approximations in some detail.
+Example 6.1.
+Consider the linear ansatz
+(74) $\tilde{\beta}(x;\eta)=Bx+d,\qquad\tilde{\Phi}(x;\alpha)=Ax+c$
+for suitable matrices $A,B$ and vector $c,d$; i.e., $\eta=(d,B)$ and $\alpha=(c,A)$. Taking variations with respect to these parameters in (72) leads to
+(75) $A_{t}=C_{t}^{px}(C_{t}^{xx})^{-1},\qquad c_{t}=\mu^{p}_{t}-A_{t}\mu^{x}_{t}$
+and
+(76) $B_{t}=-\frac{1}{2}\Sigma(C_{t}^{xx})^{-1},\qquad d_{t}=-B_{t}\mu_{t}^{x},$
+respectively.
+Here $C_{t}^{px}$ and $C_{t}^{xx}$ denote the empirical covariance matrices between $\{P_{t}^{(i)}\}$ and $\{X_{t}^{(i)}\}$, and $\{X_{t}^{(i)}\}$ with itself, respectively.
+The empirical mean of $\{X_{t}^{(i)}\}$ and $\{P_{t}^{(i)}\}$, respectively, is denoted by $\mu_{t}^{x}$ and $\mu_{t}^{p}$, respectively.
+Upon substituting these parameters into (72), the following Hamiltonian equations of motion arise from (63):
+(77a) $\dot{X}_{t}^{(i)}$ $=b(X_{t}^{(i)})-G(X_{t}^{(i)})RG(X_{t}^{(i)})^{\rm T}P_{t}^{(i)}+\frac{1}{2}\Sigma(C_{t}^{xx})^{-1}(X_{t}^{(i)}-\mu_{t}^{x}),$
+(77b) $-\dot{P}_{t}^{(i)}$ $=(Db(X_{t}^{(i)}))^{\rm T}P_{t}^{(i)}+\nabla_{x}c(X_{t}^{(i)})-\frac{1}{2}\nabla_{x}\|RG(X_{t}^{(i)})^{\rm T}P_{t}^{(i)}\|^{2}_{R}\,+$
+(77c) $\qquad\qquad\frac{1}{2}\Sigma(C_{t}^{xx})^{-1}\left(P_{t}^{(i)}-\mu_{t}^{p}-C^{px}_{t}(C_{t}^{xx})^{-1}(X_{t}^{(i)}-\mu_{t}^{x})\right)\,-$
+(77d) $\qquad\qquad\frac{1}{2}C_{t}^{px}(C_{t}^{xx})^{-1}\Sigma(C_{t}^{xx})^{-1}(X_{t}^{(i)}-\mu_{t}^{x}).$
+Note that derivatives of the parameters with respect to the particle positions and momenta do not appear because of (73c)–(73d).
+Such a linear approximation of the McKean–Pontryagin minimum principle is of interest for receding horizon stochastic optimal control problems as they arise from NMPC.
+See Section 7.3 for a numerical illustration.
+We close this section by mentioning that the deep learning approaches exploited in [8, 9, 25, 15] can also be extended to our McKean–Pontryagin formulations.
+In this case, one would utilize global parametric ansatz functions $\beta_{t}(x)=\tilde{\beta}_{t}(x;\eta)$ and $\nabla_{x}\phi_{t}(x)=\tilde{\Phi}_{t}(x;\alpha)$, $t\in[0,T]$, and directly substitute them into the action principle (21).
 7. Numerical experiments
-We present results from two numerical examples.
-Both cases involve infinite horizon cost functions of the form (3) and controlled ODEs.
-We employ the initial value formulation (44) and its numerical approximation using discrete Schrödinger bridges as outlined in Section 5.2.
+We present results from three numerical experiments.
+The first two experiments involve infinite horizon cost functions of the form (3) and controlled ODEs.
+We employ the initial value formulation (47) with the advection term in (47b) approximated using Nadaraya–Watson kernel regression [3]; i.e.,
+(78) $\nabla_{x}\phi_{t}(x)\approx\breve{y}_{t}(x):=\frac{\sum_{i=1}^{M}\exp\left(-\frac{1}{2\delta}\|x-X_{t}^{(i)}\|^{2}\right)P_{t}^{(i)}}{\sum_{i=1}^{M}\exp\left(-\frac{1}{2\delta}\|x-X_{t}^{(i)}\|^{2}\right)}$
+for suitable parameter $\delta>0$ and
+(79) $D_{x}^{2}\phi_{t}(X_{t})\dot{X}_{t}\approx\frac{1}{\Delta t}\left(\breve{y}_{t}(X_{t}+\Delta t\dot{X}_{t})-\breve{y}_{t}(X_{t})\right).$
+The operator $\mathcal{L}_{\beta}$ is approximated numerically using discrete Schrödinger bridges as outlined in Section 6.2.
 In both cases, we add a diffusion (viscosity) regularization in (22) in line with the concept of viscosity solutions to the associated HJB equations and the vanishing viscosity method [6].
-6.1. Inverted pendulum
+The final experiment returns to the inverted pendulum from the first experiment but now in the context of NMPC and receding horizon optimal control.
+Here we rely on an implementation of formulation (77).
+7.1. Inverted pendulum: Infinite horizon control
 Figure 1.
@@
 (81) $c(x)=\frac{\alpha}{2}\left(\theta^{2}+v^{2}\right)$
-with $\alpha=25$ and set $R=1$ in (4). We use the discount factor $\gamma=1$ in (3). Note that $(0,0)$ is an unstable equilibrium point under $U_{t}\equiv 0$.
-In order to implement and, in fact, regularize the mean-field PMP, we add diffusion (viscosity) by setting $\Sigma=0.1I$ in (22).
-We initialize the particles at the stable equilibrium point $x_{\rm s}=(\pi,0)$ with added Gaussian noise of variance $0.1I$; i.e.
-$X_{0}^{(i)}\sim{\rm N}(x_{\rm s},0.1I)$.
-The initial momenta are set to $P_{0}^{(i)}=-2SX_{0}^{(i)}$ with
-(64) $S=\left(\begin{array}[]{cc}0&1\\ 1&0\end{array}\right).$
-The discrete Schrödinger bridge based formulation from Section 5.2 has been implemented with ensemble size $M=200$ and $\varepsilon=\Delta t$.
-The resulting evolution equations (44) have been simulated over 8000 Euler time-steps with time-step $\Delta t=0.025$.
+with $\alpha=30$ and set $R=1$ in (4). We use the discount factor $\gamma=1.5$ in (3). Note that $(0,0)^{\rm T}$ is an unstable equilibrium point under $U_{t}\equiv 0$.
+[…]
+The discrete Schrödinger bridge based formulation from Section 6.2 has been implemented with ensemble size $M=200$ and $\varepsilon=2\Delta t$.
+The resulting evolution equations (47) have been simulated over 8000 Euler time-steps with time-step $\Delta t=0.05$.
 The results can be found in Figure 1.
@@
 The resulting final control terms
-[…]
+(82) $U_{t}^{(i)}=-G(\theta_{t}^{(i)})^{\rm T}P_{t}^{(i)},\qquad G(\theta)=\left(\begin{array}[]{c}0\\ \cos(\theta)\end{array}\right),$
 strongly depend on $\theta$ while the dependence on $v$ is less structured.
@@
 Hence a balance of computational cost and state space exploration via added diffusion is required in practice and needs to be explored further.
-6.2. Lorenz-63
+7.2. Lorenz-63: Infinite horizon control
 Figure 2.
-Controlled Lorenz-63 model: Displayed is the ${\rm x}$-component of a single particle trajectory chosen at random from the ensemble $\{X_{t}^{(i)}\}$ and its associated control $U_{t}^{(i)}$ as a function of time.
+Controlled Lorenz-63 model: Displayed is the three-dimensional trajectory of the particle $\{X_{t}^{(i)}\}$ with $i=1$ over the time interval $t\in[0,100]$.
+After an initial transient, the trajectory enters a quasi-period orbit.
 We follow [18] and consider an optimal control problem for the Lorenz-63 model
-(66) $\dot{X}_{t}=f(X_{t})+U_{t},\qquad f(X)=\left(\begin{array}[]{l}\sigma({\rm y}-{\rm x})\\ -{\rm x}{\rm z}+r{\rm x}-{\rm y}\\ {\rm x}{\rm y}-b{\rm z}\end{array}\right),$
+(83) $\dot{X}_{t}=b(X_{t})+GU_{t},\qquad b(x)=\left(\begin{array}[]{l}\sigma({\rm y}-{\rm x})\\ -{\rm x}{\rm z}+r{\rm x}-{\rm y}\\ {\rm x}{\rm y}-b{\rm z}\end{array}\right),$
 in the state variable $x=({\rm x},{\rm y},{\rm z})^{\rm T}\in\mathbb{R}^{3}$ and with parameters $\sigma=10$, $r=28$, and $b=8/3$ [19].
+The scalar-valued control $U_{t}$ acts only on the ${\rm x}$-component; i.e.
+$G=(1,0,0)^{\rm T}$.
 The control problem is to restrict solutions of (83) to ${\rm x}_{t}\geq 0$ and the authors of [18] introduce the running cost
@@
 Contrary to the setting of [18], we have implemented a discounted cost function (3) with $\gamma=1$ instead of a receding horizon NMPC approach [11, 30].
-We have implemented the Schrödinger bridge based mean-field PMP formulation with $\varepsilon=0.1$, step-size $\Delta t=0.001$, and $M=100$ particles.
+We have implemented the Schrödinger bridge based McKean–Pontryagin formulation with $\varepsilon=0.1$, step-size $\Delta t=0.001$, and $M=100$ particles.
 We again added artificial diffusion (viscosity) via $\Sigma=0.5I$ in (22).
-The results can be found in Figures 2 and 3.
+The parameter $\delta$ in (78) is set to $\delta=1$.
+The initial momenta are $P_{0}^{(i)}=0$.
+The results can be found in Figures 3 and 2.
 The impact of the control can be clearly detected.
@@
 The initial spiral towards to quasi-periodic attractor is due to the chosen initialization
-(68) $X_{0}=(7.8590;7.1136;27.2293)^{\rm T}+\Xi_{0},\qquad\Xi_{0}\sim{\rm N}(0,0.1I),$
+(85) $X_{0}^{(i)}=(7.8590,7.1136,27.2293)^{\rm T}+\Xi_{0}^{(i)},\qquad\Xi_{0}^{(i)}\sim{\rm N}(0,0.1I),$
 of the Lorenz-63 problem. Our results are in qualitative agreement with the findings from [18] which are based on a combination of data assimilation [33, 10] and NMPC [11, 30].
-Figure 3. Controlled Lorenz-63 model: Displayed is the three-dimensional trajectory of a single particle chosen at random from the controlled ensemble $\{X_{t}^{(i)}\}$.
+Figure 3.
+Controlled Lorenz-63 model: Displayed is the dependence of the computed controls $U_{t_{\rm end}}^{(i)}$ as a function of the three components of $X_{t_{\rm end}}^{(i)}$ at final time $t_{\rm end}=100$ for $i=1,\ldots,M$, $M=100$.
+7.3. Inverted pendulum: Receding horizon control
+We return to the inverted pendulum formulation (80) from Section 7.1 but now in the receding horizon setting of NMPC.
+A control is computed over receding windows of length $T=0.2$ by implementing (77) with running cost given by (81) with $\alpha=30$ and terminal cost
+(86) $f(x)=\frac{\alpha_{T}}{2}\theta^{2}$
+with $\alpha_{T}=200$. The computed closed loop control of the form
+(87) $u_{t}(x)=-G(\theta)^{\rm T}(A_{t}x+c_{t});$
+compare (82) and (75), is then applied to (80) over a prediction interval of length $\Delta\tau=0.02$ with diffusion (Brownian motion) added and $\Sigma=0.01I$.
+All equations are integrated numerically with step-size $\Delta t=0.002$ and ensemble size $M=100$.
+The initial ensemble is drawn from the Gaussian ${\rm N}((1,0)^{\rm T},0.01I)$.
+We initialize each solution of (77) with the positions from the last prediction step and the momenta from the previous optimization horizon and then solve (77) forward and backward once over a time window of length $T=0.2$.
+Numerical results can be found in Figure 4.
+We find that the receding horizon control formulation of the McKean–Pontryagin formulation is able to stabilize the unstable equilibrium $(0,0)^{\rm T}$.
+The remaining fluctuations in the ensemble are due to the non-vanishing diffusion which guarantee an exploration of state space in the vicinity of $(0,0)^{\rm T}$.
+Again an optimal choice of $\Sigma$ will depend on the particular problem formulation.
+It should be noted that the variational formulation (77) leads to local and linear approximations of the control while the approach used in Section 7.1 delivers globally valid and nonlinear controls for the inverted pendulum.
+At the same time, the local approximation is easier to compute and suits NMPC.
+Figure 4.
+NMPC for inverted pendulum: Displayed are the time evolution of the ensemble of angles and velocities as a function of time.
+Starting from an ensemble centered at $(1,0)^{\rm T}$, the controlled ensemble eventually samples the vicinity of the unstable equilibrium $(0,0)^{\rm T}$.
+The magnitude of the fluctuations depends on the magnitude of the added diffusivity.
+Those fluctuations decrease as $\Sigma\to 0$.
 8. Conclusions
 A deterministic mean-field formulation of the PMP has been presented.
-Similar to the well-known forward-backward SDE formulation of the stochastic PMP [2] and its numerical implementation; see e.g.
-[5], a regression problem (27) arises when discretized as an interacting particle system.
+Similar to the well-known forward-backward SDE formulation of the stochastic PMP [5] and its numerical implementation, a regression problem (27) arises when discretized as an interacting particle system.
 Several numerical implementations have been discussed.
-Contrary to its stochastic counterpart, the mean-field approach leads to smooth particle trajectories which should help with reducing the variance in the involved approximations.
-Furthermore, the freedom in the choice of the gauge variable $\beta_{t}$ allows for a decoupling of the forward and backward ODEs in $X_{t}$ and $P_{t}$, respectively.
+Contrary to its stochastic counterpart, the McKean–Pontryagin minimum principle leads to smooth particle trajectories which should help with reducing the variance in the involved approximations.
+The freedom in the choice of the gauge variable $\beta_{t}$ allows for a decoupling of the forward and backward ODEs in $X_{t}$ and $P_{t}$, respectively.
 Furthermore, a forward simulation of the coupled Hamiltonian ODEs is possible in case of infinite horizon optimal control problems.
 Numerical results have been presented for a controlled inverted pendulum as well as a controlled Lorenz-63 system.
-While the numerical experiments focus on infinite horizon control problems and the Schrödinger bridge approximation, application of the proposed methodology to NMPC [23, 7] and variational approximations will be investigated in future work.
-Another area of interest is provided by extensions to partially observed optimal control problems [22, 1, 17, 29, 24].
+An application of the proposed methodology to NMPC [30, 11] and a variational approximation of the McKean–Pontryagin formulation has also been investigated numerically.
+An important area of future work is provided by extensions to partially observed optimal control problems [29, 1, 23, 36, 31].
+Here a combination of the control formulation (77) with the ensemble Kalman filter [4] appears as a particularly promising direction in the context of NMPC.
 Acknowledgment. This work has been funded by Deutsche Forschungsgemeinschaft (DFG) - Project-ID 318763901 - SFB1294.